Information Sources: the Internet and Computer-Mediated Communication

John December (decemj@rpi.edu)

March 11, 1994

Release: 11 Mar 94; Release 3.30

Copyright \begin{matrix}c \\ ◯ \end{matrix}1994 John December (decemj@rpi.edu). You may use this document for any personal or edu-

cational purpose. For-profit distribution requires my permission. DISCLAIMER: Provided “as is” without expressed or implied warranty. PURPOSE: to list pointers to information describing the Internet, computer networks, and issues related

to computer-mediated communication (CMC). Topics include the technical, social, cognitive, and psychological aspects.

AUDIENCE: those getting started in understanding the Internet and CMC; for those experienced, it

compactly summarizes sources of information. ASSUMPTION: you have access to and know how to use finger, ftp, telnet, email, or Usenet newsgroups. NOTES:

1. Respect access privileges. (See Net Etiquette Guide, below.)

2. This information changes quite a bit; Additions/comments welcome.

3. Use Archie to find a file if it is not at the site given here.

4. Learn to use Gopher, Veronica, WAIS, WWW, to find more information.

Contents

1 HOW TO USE THIS DOCUMENT 2

2 THE INTERNET AND SERVICES 3

3 INFORMATION SERVICES/ELECTRONIC PUBLICATIONS 19

4 SOCIETIES AND ORGANIZATIONS 23

5 NEWSGROUPS 25

6 SELECTED BIBLIOGRAPHY 26

7 DESCRIPTION OF ITEMS 30

1 HOW TO USE THIS DOCUMENT 2

1 HOW TO USE THIS DOCUMENT

This document is formatted for LaTEX(for conversion to Postscript.) See the readme file for information on different formats for this document.

The tables in this document present information about the Internet. Tables of network-accessible

information are in the following form:

NETWORK ACCESSIBLE INFORMATION

ITEM NAME AM

PARAMETERS PARAMETERS

Where

• ITEM NAME = a short name describing the information or service.

• AM (ACCESS METHOD) = one of email, finger, ftp, gopher, http, phone, telnet, unix, Usenet news, wais.

• PARAMETERS = further access information depending on ACCESS METHOD:

ACCESS METHOD AND PARAMETERS

ACCESS METHOD PARAMETERS PARAMETERS
email  address  “letter-body”
finger  address  port #
ftp  host  path-to-directory/file-name
gopher  host
http host path-to-directory/file-name
news  newsgroup-name
phone  phone #
telnet host port # ‘login-name’
unix  command
wais  host  path

If any of the parameters are not on the line, they are not needed.

• All ftp access uses login-name ‘anonymous’ unless stated otherwise.

For information that is paper-based:

PAPER-BASED INFORMATION

Author year title

publisher or further information

INFORMATION ABOUT UPDATES OF THIS DOCUMENT

Internet-cmc information ftp ftp.rpi.edu pub/communications/internet-cmc .readme

Internet-cmc tips ftp ftp.rpi.edu pub/communications/internet-cmc .use

2 THE INTERNET AND SERVICES 3

2 THE INTERNET AND SERVICES

This section lists information about the Internet, services available on it, and topics related to computer networking.

INTERNET DESCRIPTIONS NEW USER INTRODUCTION/MOTIVATION

Gold in Networks! ftp

nic.merit.edu documents/fyi/fyi 10.txt

Hitchikers Guide ftp

nic.merit.edu documents/rfc/rfc1118.txt

Internet Index ftp crl.dec.com

pub/misc/internet-index.txt

New User’s Questions ftp

nic.merit.edu documents/fyi/fyi 04.txt

Surfing the Internet ftp nysernet.org pub/resources/guides/surfing.2.

0.3.txt

What is the Internet? ftp

nic.merit.edu documents/fyi/fyi 20.txt

Internet Services FAQ ftp rtfm.mit.edu pub/usenet/news.answers/interne

t-services/faq

COMPREHENSIVE GUIDES

AARNet Guide ftp

aarnet.edu.au pub/resource-guide/

AARNet User Guide ftp

aarnet.edu.au pub/user-guide/

Big Dummy’s Guide ftp ftp.eff.org

pub/Net info/Big Dummy/

Big Dummy Web http www.eff.org

papers/bdgtti/bdgtti.html

Big Dummy Search http alpha.acast.nova.edu cgi-bin/srch.cgi/search/bigdumm

y/mylist

CERFNet Guide ftp

nic.cerf.net cerfnet/cerfnet info/

Desktop Internet ftp ftp.uwp.edu

pub/msdos/dir/

De Presno Gopher gopher digital.cosn.org 11/Networking%20Information/Ref

erence/The%20Online%20World%20- %20Odd%20de%20Presno

De Presno Guide ftp ftp.eunet.no

pub/text/

De Presno Guide Email mail

LISTSERV@vm1.nodak.edu “get tow where”

DDN New User Guide ftp nic.ddn.mil

netinfo/nug.doc

InfoPop ftp ftp.gmu.edu library/

Internet Companion(parts) ftp

world.std.com OBS/The.Internet.Companion/

Internet Guide ftp sunsite.unc.edu pub/docs/about-the-net/libsoft/

guide1.txt

Meng’s http

ccat.sas.upenn.edu mengwong/guide.html

Neophyte ftp hydra.uwo.ca pub/libsoft/NETWORK KNOWLEDGE f

or the NEOPH.TXT

NSF Resource Guide ftp

ds.internic.net resource-guide/overview

NWNet Internet Guide ftp

ftphost.nwnet.net user-docs/nusirg/README.nusirg

NYSERNet Internet Guide ftp

nysernet.org pub/guides/Guide.V.2.2.text

SURANet Internet Guide ftp ftp.sura.net

pub/nic/

SURFnet Guide ftp

ftp.nic.surfnet.nl surfnet/user-support/docs/train ing/

Zen/Art of Internet ftp csn.org

pub/net/zen/

SPECIALIZED GUIDES OF GENERAL INTEREST

Agricultural Guide ftp sunsite.unc.edu pub/docs/about-the-net/libsoft/

agguide.dos

Electric Mystics Guide ftp

panda1.uottawa.ca pub/religion/

Library Resources ftp

dla.ucop.edu pub/internet/libcat-guide

BUBL gopher

bubl.bath.ac.uk

BUBL Web http bubl.bath.ac.uk BUBLHOME.html

EXPLORING

DOS Internet Kit ftp tbone.biol.scarolina.edu

pub/kit/

DOS Internet Kit Web http tbone.biol.scarolina.edu

dean2/kit/kit.html

Internet Beginner ftp

oak.oakland.edu pub/msdos/info/bgi13.zip

Internet Cruise ftp nic.merit.edu

resources/

Internet Explorer’s Kit ftp sunsite.unc.edu pub/docs/about-the-net/libsoft/

explorer.doc

Internet Hunt ftp ftp.cni.org

pub/net-guides/i-hunt/

Internet Hunt Gopher gopher

gopher.cic.net

Internet Navigating ftp ubvm.cc.buffalo.edu

navigate/

Internet Mining ftp

ftp.ucdavis.edu ucd.netdocs/mining/mining readm e

Internet Safari ftp nic.cerf.net cerfnet/cerfnet k12 programs/ce

rf n safari/safari guide/safari -guide.txt

TRAINING

Competencies mail

listserv@uhupvm1.uh.edu “GET CORBIN PRV4N6”

Discussion help mail

rre-request@weber.ucsd.edu Subject: archive send courtesy

Gopherin ftp ubvm.cc.buffalo.edu

gophern/

HTML Guide http

www.ncsa.uiuc.edu demoweb/html-primer.html

HTML Guide/Flynn http www.ucc.ie

info/net/htmldoc.html

HTML Tutorial http

fire.clarkson.edu doc/html/htut.html

HWGUIDE ftp sunsite.unc.edu pub/docs/about-the-net/libsoft/

hwguide.txt

LeJeune Course ftp

pilot.njin.net pub/Internet-course/

Milles ftp

sluaxa.slu.edu pub/millesjg/interlaw.txt

Monsarrat ftp

wilma.cs.brown.edu pub/internet course.tar.Z

Navigating ftp

ftp.sura.net pub/nic/training/

Navigating gopher

jake.esu.edu 11/Help/net stuff/training

Net workbook mail listserv@bingvmb.bitnet “GET NETKNOW NEOPHYTE BI-L”

Newcastle/Tyne ftp mailbase.ac.uk pub/lists/itti-networks/files/t

ms-list.txt

Schneider Bib ftp

alexia.lis.uiuc.edu pub/training.bib

Start ftp

sluaxa.slu.edu pub/millesjg/newusers.faq

Syracuse Courses http

eryx.syr.edu bmis/bmis.html

Trainmat gopher

trainmat.ncl.ac.uk

Trainpack ftp

ftp.ncl.ac.uk pub/network-training/README

Wheeler ftp s850.mwc.edu

nettrain/nettrain.00readme

Unix tutorials gopher

jake.esu.edu 11/Help/Tutorials

Unix Help http

alpha.acast.nova.edu Unixhelp/TOP .html

URL gopher

gopher.well.sf.ca.us 00/matrix/internet/curling.up.0 2

URL guide http

www.ncsa.uiuc.edu demoweb/url-primer.html

Usenet and trn ftp

cs1.presby.edu pub/trn-intro/

VAX/VMS Networking ftp

ftp.temple.edu pub/info/help-net/vms-mail.guid e

Windows-based ftp

WinFtp.CICA.Indiana.Edu ADMINISTRATIVE/TECHNICAL/HISTORY

Internet Growth ftp tic.com

matrix/growth/internet/

Internet History ftp

umcc.umich.edu pub/users/seraphim/doc/nethist8 .txt

Internet History ftp

cs.beloit.edu Public/Internet/History/

Internet Monthly Report ftp nic.merit.edu internet/newsletters/internet.m

onthly.report/

Internet Maps (Europe) ftp eunet.fi

nic/pub/netinfo/maps/

Internet Maps (Many) ftp ftp.uu.net

inet/maps/

Internet Maps (NSFNET) ftp nic.merit.edu

maps/

Internet Maps (SuraNet) ftp ftp.sura.net

pub/maps/

Internet Size ftp ftp.nisc.sri.com

pub/zone/

Internet Statistics ftp

nic.merit.edu nsfnet/statistics/

Web country maps http www.tue.nl maps.html

INFO REPOSITORIES COMPREHENSIVE COLLECTIONS

Merit ftp nic.merit.edu

READ.ME

Merit Gopher gopher

nic.merit.edu

NIC/DDN ftp nic.ddn.mil

NETWORK INFO CENTER

InterNic Telnet telnet ds.internic.net

‘guest’

InterNIC Ftp ftp

ds.internic.net pub/InterNIC-info/internic.info

InternNIC via email mail mailserv@internic.net

“help”

InterNIC Gopher gopher

rs.internic.net

InterNIC Home Page http

www.internic.net internic.html

InterNIC Database Services ftp

ds.internic.net

InterNIC Reg Services ftp rs.internic.net

netinfo/

InterNIC Info Sources ftp is.internic.net infosource/infosource-contents

DOCUMENT SERIES

FYI ftp

nic.merit.edu documents/fyi/

Internet Drafts ftp

nic.merit.edu documents/internet-drafts/

Internet Standards ftp

nic.merit.edu documents/std/

RFC ftp

nic.merit.edu documents/rfc/

RFC Web http www.cis.ohio-state.edu hypertext/information/rfc.html INTERNET HOME PAGES

Duke’s NetCom http

porter.netcom.duke.edu

Internet Web Text http www.rpi.edu

Internet/Guides/decemj/text.htm l

Italian Net Tools http

www.mi.cnr.it NIR-IT/NIR-map.html

GNN’s Internet http

nearnet.gnn.com wic/internet.toc.html

Nova Links http alpha.acast.nova.edu

start.html

Stanford http

med-www.stanford.edu hypertext/net/Net.html

NCIH http

www.cnidr.org welcome.html 5050

NOSC Page http

white.nosc.mil internet.html

Yale http

www.cs.yale.edu HTML/WorldWideWebTop.html

WWW World Maps http wings.buffalo.edu

world

WWW European Maps http www.tue.nl

maps.html

SOURCES OF NETWORKING, COMPUTING AND RELATED INFORMATION

ACM Gopher gopher

acm.org

ACM WWW Home Page http

info.acm.org

ALAWON archive mail

listserv@uicvm.uic.edu “send ala-wo filelist”

CMD http

www.cmd.uu.se

CNI Gopher gopher

gopher.cni.org

CNI search telnet gopher.cni.org

‘brsuser’

CPSR Gopher gopher

gopher.cpsr.org

CTR/Columbia http

www.ctr.columbia.edu CUCTR Home.html

Communications Res Cen http debra.dgbt.doc.ca

index.html

CERN http

www1.cern.ch

CWI http

buizerd.cwi.nl default.html

Dern’s Info gopher gopher.internet.com

11/ 2200

EARN Gopher gopher

gopher.earn.net

EFF Gopher gopher

gopher.eff.org

Hacking and Phreaking http www.ugcs.caltech.edu

phd

HelpNet Archives ftp

ftp.temple.edu pub/info/help-net

HCI ftp site ftp archive.cis.ohio-state.edu

pub/hcibib/

HCI Launching Pad http hydra.bgsu.edu

HCI/

HPCC Office Web Page http

www.hpcc.gov

HPCC Gopher gopher gopher.hpcc.gov

70

HPCWire telnet

hpcwire.ans.net

IIASA gopher

/gopher.iiasa.ac.at

ICSI http

http://http.icsi.berkeley.edu

IEEE Gopher gopher

info.ieee.org

IEEE Tech mail info.new.tech@ieee.org

“anything”

ISOC Gopher gopher gopher.isoc.org ISS http

www.iss.nus.sg

ISRI http

www.isri.unlv.edu

Introducing the Internet ftp

nic.merit.edu introducing.the.internet/

Library of Congress ftp ftp.loc.gov

pub/iug/

LIT http litsun.epfl.ch

80

Mem U of Newfoundland gopher cwis.ucs.mun.ca 11/Venturing%20into%20the%20Int ernet

MetaCenter http www.ncsa.uiuc.edu General/MetaCenter/MetaCenterHo

me.html

Meta Net Gopher gopher

tmn.com

MSU Net and DB gopher

burrow.cl.msu.edu

MSEN Gopher gopher

gopher.msen.com

NEARnet ftp nic.near.net

docs/

NIST Web Page http

www.nist.gov welcome.html

NIST Gopher gopher

gopher-server.nist.gov

NSF Gopher gopher stis.nsf.gov

11 70

NTTC http iridium.nttc.edu

nttc.html

NYSERnet ftp

nysernet.org pub/resources/guides/

OPTOLINK ftp

mom.spie.org

OSS gopher

gopher.oss.net

PSGnet/RAINet gopher

gopher.psg.com

RICIS http

http://rbse.jsc.nasa.gov

RIPE Gopher gopher

gopher.ripe.net

RIPE Web http

www.ripe.net

Singapore DMC http

king.ncb.gov.sg

Software Tool and Die gopher

world.std.com

SRI’s Network Info ftp sri.com

netinfo/

SURAnet ftp

ftp.sura.net pub/README

SWITCHinfo ftp

nic.switch.ch info service/

Telecomm Archives ftp lcs.mit.edu

telecom-archives/

Telecomm research ftp

ftp.ctr.columbia.edu CTR-Research/

UNC’s Sunsite ftp

sunsite.unc.edu pub/docs/about-the-net/

USENET Repository ftp rtfm.mit.edu

pub/usenet/

USENET Periodic Postings ftp rtfm.mit.edu pub/usenet/news.answers/periodi

c-postings/

UUNET archive ftp ftp.uu.net

uunet-info/

UWO’s NIC ftp julian.uwo.ca

nic/

Washington and Lee gopher

liberty.uc.wlu.edu

WATERS http

www.cs.odu.edu WATERS/WATERS-GS.html

WELL Matrix gopher gopher.well.sf.ca.us 11/matrix

SERVICES AND TOOLS INFORMATION AND SERVICES LISTS

Cyberpoet Guide ftp ftp.eff.org

pub/Net info/Cyber/cyberpoet.gv c

Fishnet http

www.cs.washington.edu homes/pauld/fishnet/

Freeside FAQ http www.seas.upenn.edu

mengwong/fsfaq.html

Gopher Jewels gopher cwis.usc.edu 11/Other Gophers and Informatio

n Resources/Gopher-Jewels

Gopher Jewels Web http

galaxy.einet.net GJ/index.html

Gopher Jewels ftp usc.edu

pub/gopher/

Gopher Jewels http

galaxy.einet.net GJ/index.html

Gopher Jewels Info mail

listproc@einet.net “get gopher-lists.txt”

Hot/Cool List http 141.214.4.176

uwi/reviews.html

Internet Handbook ftp sri.com netinfo/internet-technology-han dbook-contents

Information Sources ftp ftp.rpi.edu pub/communications/internet-cmc .readme

Internet Meta-Index http www.ncsa.uiuc.edu SDG/Software/Mosaic/MetaIndex.h tml

Internet Reviews gopher

jupiter.willamette.edu 11/library/reviews

Joel’s List http www.cen.uiuc.edu

jj9544/index.html

MaasInfo Indexes ftp ftp.unt.edu

pub/articles/maas/maasinfo.file s

May List ftp

aug3.augsburg.edu files/bbs lists/nal006.txt

Must Have List ftp pilot.njin.net pub/Internet-course/musthave-li

st.txt

Networking Overview http

web.doc.ic.ac.uk bySubject/Networking.html 80

Planet Earth http white.nosc.mil

info.html

Smith’s BigFun List ftp

owl.nstn.ns.ca pub/netinfo/bigfun.txt

Stanford WWW Links http

www.stanford.edu otherLinks.html

ThesisNet FAQ http www.seas.upenn.edu

mengwong/thesisfaq.html#ftp

Yale Overview http

www.cs.yale.edu HTML/WorldWideWebTop.html

Yanoff’s Services List ftp

csd4.csd.uwm.edu pub/inet.services.txt

Yanoff List HTML ftp csd4.csd.uwm.edu pub/inet.services.html

LISTS OF TOOLS

Best of Web http wings.buffalo.edu

contest

Internet Browsers http

life.anu.edu.au links/syslib.html

Internet Tools EARN ftp ns.ripe.net

earn/earn-resource-tool-guide.t xt

Internet Tools/EARN HTML http

www.huji.ac.il www help/earn.html

Internet Tools NIR ftp

mailbase.ac.uk pub/lists/nir/files/nir.status. report

Internet Tools/December HTML http www.rpi.edu

Internet/Guides/internet-tools. html

Internet Tools/December ftp ftp.rpi.edu

pub/communications/internet-too ls

Internet Systems UNITE ftp mailbase.ac.uk pub/lists/unite/files/systems-l ist.txt WEB SEARCHING TOOLS

Archie via Web http hoohoo.ncsa.uiuc.edu

archie.html

Gopher Jewels Search http

galaxy.einet.net gopher/gopher.html

GNA Meta-Library http

uu-nna.mit.edu uu-nna/meta-library/index.html 8001

Hyper-Index http

flick.lerc.nasa.gov cgi-bin/h-index/query-index-for m

JumpStation http www.stir.ac.uk

jsbin/js

NorthStar http comics.scs.unr.edu

top.html? 7000

SUSI http

web.nexor.co.uk susi/susi.html

WWW Catalog http cui www.unige.ch

w3catalog

WWW Search Engines http

cui www.unige.ch meta-index.html

WWW Worm http

www.cs.colorado.edu home/mcbryan/WWWW.html

WWW Wanderers http web.nexor.co.uk mak/doc/robots/robots.html

TOOLS

For a list of tools and documentation related to each tool, see the Internet Tools Summary/December, listed above.

SERVICES

Anonymous post mail anonymus+ping@tygra.michigan.co m

“help”

Anonymous post mail help@anon.penet.fi

“help”

Anonymous remail mail

remail@tamsun.tamu.edu Subject: remail help

BITFTP mail Bitftp@PUCC.Princeton.edu

“help 13-Dec-93”

CARL Systems telnet

database.carl.org

Correct Time/NBS telnet india.colorado.edu

13

ENewsstand telnet enews.com ‘enews’

ENewsstand gopher gopher

gopher.internet.com

Email FTP mail ftpmail@decwrl.dec.com

“help”

Email Gopher mail gophermail@ncc.go.jp

“help”

Email Gopher mail gophermail@calvin.edu

“help”

Email WWW mail listserv@info.cern.ch

“www URL”

Email Usenet mail hierarchy-group-name@cs.utexas. edu Subject: Your Subject “Your Co ntents”

FAXNET mail info@awa.com

“help”

Finger database finger

help@dir.su.oz.au

Hytelnet telnet telnet

Internet Marketing mail

market@internet.com “send market info””

Knowbot telnet info.cnri.reston.va.us

185

Mail name server mail dns@grasp.insa-lyon.fr

“help”

Softlock mail Info@SoftLock.COM

“send info”

WorldWindow telnet library.wustl.edu

NETWORKING

BITNET intro mail

listserv@bitnic.educom.edu “send BITNET INTRO”

CERT ftp cert.org pub/

DoD Security ftp asc.dtic.dla.mil

pub/tafim/

FidoNet Guide gopher digital.cosn.org 11/Networking%20Information/Ref

erence/The%20BIG

FidoNet News ftp

rtfm.mit.edu pub/usenet/comp.org.fidonet/

FidoNet Nodes ftp

genome.wi.mit.edu wais-sources/fidonet-nodelist.s rc

FidoNet Gateway ftp csn.org

pub/mail/internet.fidonet

Hypercard TCP/IP http ericmorgan.lib.ncsu.edu staff/morgan/papers/tcp-communi

cations/tcp-communications.html

INET 93 ftp mordor.stanford.edu

pub/inet93/

International Connect ftp

ftp.cs.wisc.edu connectivity table/

Internet Country Codes ftp rtfm.mit.edu pub/usenet/news.answers/mail/co

untry-codes

Internet Domain Names ftp

nic.merit.edu documents/rfc/rfc1394.txt

Internet Root Domain ftp ftp.rs.internic.net

domain/

Internet Security ftp

nic.merit.edu cise/gao8957.txt

InterNetwork Mail ftp

csd4.csd.uwm.edu pub/internetwork-mail-guide

Intro TCP/IP ftp

nic.merit.edu introducing.the.internet/intro. to.ip

List of FTP Sites ftp

ftp.gsfc.nasa.gov pub/ftp-list

PSGnet/RAINet info gopher

rain.psg.com

NREN Information ftp

nic.merit.edu nren/INDEX.nren

NREN Recompetition ftp

nic.merit.edu cise/recompete/INDEX.recompete

OneNet phone 415-948-4775

SDSC Appl Net Res Group ftp

ftp.sdsc.edu pub/sdsc/anr/README

Sprintlink ftp

ftp.sprintlink.net

Sprintlink Gopher gopher

ftp.sprintlink.net

USENET Maps ftp gatekeeper.dec.com

pub/maps/

WWIV gopher gopher.technet.sg 11/pub/msdos/wwiv 2100 ACCESSING NETWORKS

BBS Internet List ftp sunsite.unc.edu pub/docs/about-the-net/libsoft/

internet bbs.txt

Community Networks ftp netcom.com pub/amcgee/community/communet.m sg

Connecting to Internet ftp

nic.merit.edu documents/fyi/fyi 16.txt

DLIST mail

dlist@ora.com “Please send DLIST”

Freenet papers ftp

alfred.carleton.ca pub/freenet/working.papers/

FSLIST ftp

freedom.nmsu.edu pub/docs/fslist/

Global Networks ftp

dhvx20.csudh.edu global net.dir/readme.txt

Internet Access Guide ftp

nic.merit.edu introducing.the.internet/access .guide

Internet Access ftp

sluaxa.slu.edu pub/millesjg/internet.access

Internet Modem ftp

dla.ucop.edu pub/internet/dial-access

Internet Providers ftp sri.com netinfo/internet-access-provide rs-alphabetical-listing.txt

Internet Providers Non-US ftp sri.com netinfo/internet-access-provide rs-non-us.txt

K12 Access ftp ftp.cc.berkeley.edu

k12/README

K12 Gopher gopher

gopher.cic.net 11/cicnet-gophers/k12-gopher

NII http sunsite.unc.edu

nii

Network Startup ftp ftp.psg.com

README

PDIAL ftp

rtfm.mit.edu pub/usenet/news.answers/pdial

NIXPUB ftp rtfm.mit.edu pub/usenet/alt.bbs/Nixpub Posti

ng (Long)

Registering on the Net ftp

nic.merit.edu documents/rfc/rfc1400.txt

Rural Nets mail

rjacot@cic.net

Rural Datafication gopher gopher

gopher.cic.net 11/cicnet-gophers/ruraldata-pro ject

Rural Datafication Web http www.cic.net

rd-home.html

SEA Community Network ftp atlas.ce.washington.edu pub/seattle-community-network/

SOME COMMERCIAL SERVICE (Online, Internet, Public-access) PROVIDERS

America Online phone

800-827-6364

BIX  phone 800-695-4775
CONNECT  phone 408-973-0110
Compuserve  phone 800-848-8990
Delphi  phone 800-695-4005
Dialog  phone 800-334-2564
Dow Jones  phone 800-522-3567
EUnet  http  www.eu.net
GEnie  phone 800-638-9636
MCI Mail  phone 800-444-6245

Netcom mail

info@netcom.com

PC-LINK  phone 800-827-8532
PSI  phone 703-620-6651

PSI mail

all-info@psi.com

Pipline mail infobot@pipeline.com

“help”

Prodigy phone 800-766-3449

TMN mail

info@tmn.com

The WELL mail

support@well.sf.ca.us

The World gopher

world.std.com

SOME FREE-NETS (tm) / COMMUNITY NETWORKS

Free-Net List/Gopher gopher

gopher.tamu.edu 11/.dir/freenet.dir

Free-Net List/Hytelnet http

www.cc.ukans.edu hytelnet html/FRE000.html

Blacksburg http

gopher.vt.edu/bev/index.html

CIAO! telnet

ciao.trail.bc.ca

Chebucto FreeNet http

cfn.cs.dal.ca cfn/Home.html

Cleveland telnet freenet-in-a.cwru.edu

‘new’

Digital City telnet

dds.hacktic.nl

M-Net telnet m-net.ann-arbor.mi.us

‘newuser’

Nyx telnet nyx.cs.du.edu

‘new’

NCF telnet freenet.carleton.ca

‘guest’

NS*CAN http

www.cs.dal.ca mcnp/Home.html

Toledo telnet fnet.cc.utoledo.edu

‘visitor’

Triangle NC gopher

tfnet.ils.unc.edu

Victoria BC gopher freenet.victoria.bc.ca

vifa.html

Youngstown  telnet yfn.ysu.edu  ‘visitor’

INFORMATION RETRIEVAL AND DISSEMINATION

Cataloging Internet Res mail

listserv@uhupvm1.uh.edu “get caplan prv4n2”

CHAT ftp debra.dgbt.doc.ca

pub/chat/

CHAT telnet

debra.doc.ca

CIIR http

ciir.cs.umass.edu

CNI ftp ftp.cni.org CNI/

Doc Center http

doccenter.com doccenter/home.html

DIMUND gopher

dimund.umd.edu

DIMUND FTP ftp dimund.umd.edu

pub/

CNI TopNode Project ftp ftp.cni.org

CNI/projects/topnode/

CNI Gopher gopher

gopher.cni.org

CNIDR Gopher gopher

gopher.cnidr.org

CNIDR Web Page http cnidr.org

welcome.html

IAFA ftp

archive.cc.mcgill.ca pub/Network/iafa/charter

Integrated Information ftp venera.isi.edu internet-drafts/draft-ietf-iiir

-vision-01.txt

Library Special ftp

vm1.nodak.edu nnews/nnews.1993-10

NIR Archives ftp

mailbase.ac.uk pub/lists/nir/

NIR Gopher gopher

mailbase.ac.uk

OCLC Research ftp ftp.rsch.oclc.org pub/internet resources project/

report/

PCP http

pespmc1.vub.ac.be RELATED.html

Resource Transponders ftp venera.isi.edu internet-drafts/draft-ietf-iiir

-transponders-01.txt

Retrieval success ftp mailbase.ac.uk pub/lists/unite/files/internet-

stories.txt

Riddle ftp ftp.cwi.nl

pub/RIDDLE/

Uniform Resources ftp

info.cern.ch pub/ietf/url3.txt

UNITE Archive ftp

mailbase.ac.uk pub/lists/unite/

URN ftp venera.isi.edu internet-drafts/draft-ietf-uri- resource-

names-01.txt

URI Archive ftp

archives.cc.mcgill.ca pub/mailing-lists/uri-archive

UWI http zapruder.pds.med.umich.edu uwi/uwi-info.html

DIRECTORIES

Commercial List http

tns-www.lcs.mit.edu commerce.html

CSOIRG Home Page http http2.sils.umich.edu

lou/chome.html

Dartmouth Merged SIGL ftp dartcms1.dartmouth.edu

siglists/

Electronic Conferences ftp

ksuvxa.kent.edu library/acadlist.readme

Electronic Zines/Jrl http www.acns.nwu.edu

ezines

Electronic Journals ftp ftp.cni.org

pub/net-guides/strangelove/

Electronic Magazines ftp etext.archive.umich.edu

pub/Zines/

Electronic Resources gopher watserv2.uwaterloo.ca

11/servers

EZines List ftp netcom.com

pub/johnl/zines/e-zine-list

EZines Web ftp netcom.com

pub/johnl/zines/e-zine-list.htm l

EINet Galaxy http galaxy.einet.net

galaxy.html

Finding Lists mail

listserv@vm1.nodak.edu “get NEW-LIST wouters”

Free Databases ftp

idiom.berkeley.ca.us pub/free-databases

InterActive Yellow Pages mail tuna@netcom.com Subject: yellow pages “YOUR EM AIL

ADDRESS”

Interest Groups List ftp

ftp.nisc.sri.com netinfo/interest-groups.txt

Library Access Script ftp sonoma.edu

pub/libs.tar

Library Guide ftp

ftp.utdallas.edu pub/staff/billy/libguide

Library Catalogs Noonan ftp sunsite.unc.edu pub/docs/about-the-net/libsoft/

guide2.txt

Library Catalogs St Geo ftp

nic.cerf.net internet/resources/library cata log/

Online books http

www.cs.cmu.edu Web/books.html 8001

Online bibs http

www.cs.cmu.edu Web/bibliographies.html 8001

Online journals http

www.cs.cmu.edu Web/journals.html 8001

Online libraries http

www.cs.cmu.edu Web/e-libraries.html 8001

PAML http www.ii.uib.no

magnus/paml.html

Scholarly Societies gophers gopher

watserv2.uwaterloo.ca 11/servers/campus/scholars

Subject Guides ftp una.hh.lib.umich.edu

inetdirs/

Subject Guides gopher una.hh.lib.umich.edu

11/inetdirs

Subject Guides http www.lib.umich.edu

chhome.html

Technical Reports ftp

daneel.rdt.monash.edu.au pub/techreports/

Virtual Tourist http wings.buffalo.edu

world

WAIS sources ftp

kirk.bond.edu.au pub/Bond Uni/doc/wais/readme

WAIS Directory of Servers wais cnidr.org

directory-of-servers? 210

WWW Catalogues http www.cs.colorado.edu homes/mcbryan/public html/bb/su

mmary.html

WWW Sites http www.mit.edu afs/sipb/user/mkgray/ht/compre3 .html 8001 COMMUNICATION AUDIO

Berklee gopher

gopher.berklee.edu

CERL http

datura.cerl.uiuc.edu

Internet Sound ftp ftp.cwi.nl

pub/audio/INDEX

Internet Talk Radio ftp

sunsite.unc.edu pub/talk-radio/ITRintro.readme

Internet Multicasting FAQ mail info@radio.com

“send FAQ”

Internet Multicasting WWW http

www.cmf.nrl.navy.mil radio/radio.html

Internet Talk Radio sites mail

sites@radio.com “send SITES”

Leeds/Music http

www.leeds.ac.uk music/Man/c front.html

MIDI http

www.eeb.ele.tue.nl midi/index.html

Multicast Backbone ftp

venera.isi.edu mbone/faq.txt

Mbone FAQ Web http

www.research.att.com mbone-faq.html

Music Resources http

www.music.indiana.edu misc/music resources.html

Say http utis179.cs.utwente.nl

say/form/ 8001

IUMA http

sunsite.unc.edu ianc/index.html HYPER/VIRTUAL/MULTIMEDIA

ANIMA http wimsey.com

anima/ANIMAhome.html

CGU (Manchester, UK) http

info.mcc.ac.uk CGU/CGU-research.html

File formats ftp

ftp.ncsa.uiuc.edu misc/file.formats/

GA Tech Graphics http

www.gatech.edu gvu/gvutop.html

HTML guide http www.ucc.ie

info/net/htmldoc.html

HTML guide/Tilton http

www.willamette.edu html-composition/strict-html.ht ml

HTML info http info.cern.ch hypertext/WWW/MarkUp/MarkUp.htm

l

HTML+ DTD ftp nic.merit.edu documents/internet-drafts/draft

-raggett-www-html-00.txt

Hypermedia/Internet http

life.anu.edu.au education/hypermedia.html

Hypertext terms http

info.cern.ch hypertext/WWW/Terms.html

MediaMOO telnet purple-crayon.media.mit.edu

8888

Meta VE http www.gatech.edu gvu/people/Masters/Rob.Kooper/M

eta.VR.html

Multimedia Index http

cui www.unige.ch OSG/MultimediaInfo

Multimedia Survey ftp

ftp.ed.ac.uk pub/mmsurvey/mmsurvey.txt

Multimedia Lab BU http

spiderman.bu.edu

MIT Media Lab ftp media-lab.media.mit.edu

access/

MIT Meida Lab–how to http cs.indiana.edu docproject/mit.research.how.to/

mit.research.how.to.html

MIT Telemedia http

tns-www.lcs.mit.edu tns-www-home.html

Scientific Visualization http www.nas.nasa.gov RNR/Visualization/annotatedURLs .html

Sunsite Multimedia http

sunsite.unc.edu exhibits/exex.html

TX A+M Media Lab ftp

bush.cs.tamu.edu pub/home.html

Video http

tns-www.lcs.mit.edu cgi-bin/vs/vsbrowser

VR Page/Chris Hand http

www.cms.dmu.ac.uk People/cph/vrstuff.html 9999

VR Web http guinan.gsfc.nasa.gov

W3/VR.html

VR Archive ftp sunsite.unc.edu pub/academic/computer-science/v

irtual-reality/

VSR http nfhsg3.rus.uni-stuttgart.de virtual/index.html MASS COMMUNICATION

CBC Home Page http

debra.dgbt.doc.ca cbc/cbc.html

CBC Gopher gopher

debra.dgbt.doc.ca

ClariNet mail

info@clarinet.com

College Newspapers gopher

blick.journ.latech.edu

Cyberspace Report http www.ics.uci.edu

ejw/csr/cyber.html

Dateline NBC mail

dateline@nbc.ge.com

Ham Radio http

www.acs.ncsu.edu HamRadio 80

ITR http

www.ncsa.uiuc.edu radio/radio.html

Live TV http

tns-www.lcs.mit.edu cgi-bin/vs/vsdemo

Media BBS mail

listserv@ulkyvm.louisville.edu “GET MEDIA BBSLIST”

Movies and TV http alpha.acast.nova.edu movies.html

MTV  http  mtv.com

NBC News mail

nightly@nbc.com

NPR Weekend Ed mail

wesun@cap.gwu.edu “Please send info.”

NPR Talk/Nation mail

totn@aol.com “Please send info.”

NPR Science Friday mail

scifri@aol.com “Please send info.”

Palo Alto Weekly http

www.service.com PAW/home.html

PWN mail

pwn-info@igc.apc.org

Journalism ftp rtfm.mit.edu pub/usenet/news.answers/journal

ism-net-resources

Satellite http

itre.uncecs.edu misc/sat.html

TV Schedules http

white.nosc.mil television.html

VOA ftp ftp.voa.gov

GROUP COMMUNICATION

CCCC http it.njit.edu njIT/Department/CCCC/default.ht ml

Communication Archive ftp

sunsite.unc.edu pub/academic/communications/

Communication Archive Web http sunsite.unc.edu pub/academic/communications/com

munications.html

Computer Network Conf ftp

nic.merit.edu documents/rfc/rfc1324.txt

Hypermedia/Internet http

life.anu.edu.au education/hypermedia.html

Interactive Geo http www.hcc.hawaii.edu

htbin/plotd

Internet Relay Chat(IRC) ftp cs.bu.edu

irc/support/

IRC FAQ http www.kei.com irc.html

Free for All http south.ncsa.uiuc.edu

Free.html

Multiple User Dialogue (MUD) ftp

ftp.math.okstate.edu pub/muds/misc/mud-faq/

MUD Lists ftp caisr2.caisr.cwru.edu

pub/mud/

MUD Page http math.okstate.edu

mud.html 8001

LISTSERV managing mail

listserv@uhupvm1.uh.edu “get kovacs prv2n1”

LISTSERV searching mail

listserv@ulkyvm.bitnet “get database search”

LISTSERV tips mail

listserv@bitnic.bitnet “get listserv tips” ORGANIZATIONAL COMMUNICATION

Campus Net ftp

gandalf.iat.unc.edu technote/teknote4.txt

Campus Net Bib ftp gandalf.iat.unc.edu guides/irg-15.txt

EMAIL

Email Services ftp sunsite.unc.edu pub/docs/about-the-net/libsoft/

email services.txt

Email Understanding ftp

ftp.cso.uiuc.edu doc/net/uiucnet/vol3no2.txt

Email 101 ftp

mrcnext.cso.uiuc.edu etext/etext93/email025.txt

College Email addresses ftp

rtfm.mit.edu pub/usenet/soc.college/

Finding Email addresses ftp sunsite.unc.edu pub/docs/about-the-net/libsoft/

email address.txt

Mail2Html http

neptune.corp.harris.com mail2html.html

MetaMail ftp

thumper.bellcore.com pub/nsb/README

MIME ftp

nic.merit.edu documents/rfc/rfc1341.txt

MIME overview ftp

thumper.bellcore.com pub/nsb/MIME-overview.txt

MIME info ftp ftp.uu.net

networking/mail/mime/

Pine email ftp

ftp.cac.washington.edu pine/pine.blurb

PGP Mail ftp ftp.uu.net

pub/security/pgp/

PGP/PEM http

hoohoo.ncsa.uiuc.edu docs/PEMPGP.html

PEM ftp ftp.tis.com

pub/PEM/FAQ

RIPEM http cs.indiana.edu ripem/dir.html

LANGUAGE/CULTURE/COMMUNITY/SOCIETY

Activism ftp netcom.com

pub/amcgee/activism

Anonymity FAQ ftp rtfm.mit.edu pub/usenet/news.answers/net-ano

nymity/

African ftp netcom.com

pub/amcgee/african

APC ftp igc.apc.org

pub/orgs-on-igc

Art/images gopher cs4sun.cs.ttu.edu/11/Art%20and% 20Images

Artwork ftp

sunsite.unc.edu pub/multimedia/pictures/OTIS/

ASCII Art ftp

genesis.mcs.com mcsnet.users/jorn/asciifaq.us

ASCII Art collection ftp

ftp.cs.ttu.edu pub/asciiart/

BABEL 94 ftp

ftp.temple.edu pub/info/help-net/babel94a.txt

Baylor etexts ftp ftp.byu.edu

pub/next/Literature/

Bordeaux and Prague http

mailbox.cdtl.umn.edu

CIS-AH http

web.cal.msu.edu

CTI mail

ctitext@vax.ox.ac.uk

Code of the Geeks http www.cs.odu.edu

mark/geek.html

Coke Machines http www.cs.cmu.edu afs/cs.cmu.edu/user/bsy/www/cok

e.html 8001

Community Nets Surveys http

www.cs.washington.edu research/community-networks/ 80

Community Nets ftp ftp.apple.com

alug/

Community Nets gopher gopher.well.sf.ca.us 11s/Community/communets/net.com

70

Computer Jargon search http web.cnam.fr bin.html/By Searchable Index?Ja rgon File.html

Computer Jargon ftp

aeneas.mit.edu pub/gnu/jargon-README

CMC Glossary gopher

sjumusic.stjohns.edu 11/%40uni%3acmc.glossary 1070

Computer Underground ftp ftp.eff.org pub/Publications/CuD/Papers/mey er

Computing Dictionary gopher

wombat.doc.ic.ac.uk

Computing Dictionary http

wombat.doc.ic.ac.uk

Coombs Papers gopher

coombs.anu.edu.au

CPET ftp

guvax.georgetown.edu cpet projects in electronic tex t/

Culture/Tech http english-server.hss.cmu.edu

Cyber.html

CyberCafe http

cybercafe.demon.co.uk

Cyber Papers ftp ftp.eff.org

pub/Net info/Cyber/

Cyberspace/Law ftp ftp.eff.org pub/Publications/CuD/Papers/cyb erspace

Cyberpunk FAQ ftp rtfm.mit.edu pub/usenet/news.answers/cyberpu

nk-faq

Cyberspace/Language http nearnet.gnn.com gnn/news/archives/94.01.31/MLA.

html

Encyclopedia Britannica ftp eb.com

pub/

EStyle ftp aultnis.rutgers.edu

litext

Electronic Text archive gopher fir.cic.net

00/0-README

Electronic Text ftp

guvax.georgetown.edu cpet projects in electronic tex t/

ETC http

www.lib.virginia.edu etext/ETC.html

Electronic Word ftp

press-gopher.uchicago.edu pub/Excerpts/lanham.txt

English Server http

english-server.hss.cmu.edu

FineArt Forum ftp

ra.msstate.edu pub/archives/fineart online

Friends + Partners http

solar.rtd.utk.edu friends/home.html

FreeNet93 ftp

alfred.carleton.ca pub/freenet/93conference/

Future Culture mail

future-request@nyx.cs.du.edu Subject: send faq

Gender ftp netcom.com

pub/amcgee/gender/

Gender Issues ftp alfred.carleton.ca pub/freenet/93conference/leslie

regan shade.txt

Gutenberg Project Texts ftp quake.think.com”

pub/etext/

Gutenberg Web Page http med-amsa.bu.edu Gutenberg/Welcome.html Hacker Crackdown http

www.scrg.cs.tcd.ie scrg/u/bos/hacker/hacker.html

Hacker’s Dictionary http iicm.tu-graz.ac.at

Cjargon

Humanities mail

rre-request@weber.ucsd.edu Subject: archive send humanitie s

Humanities/computing http

nearnet.gnn.com wic/hum.05.html

Hypertext/rhetoric http

fire.clarkson.edu horn/proposal-mla.html

Indigenous ftp netcom.com

pub/amcgee/indigenous/

Internet demographics ftp ftp.tic.com

survey/

Internet Glossary wais

pinus.slu.se Internet-user-glossary? 210

Internet Glossary ftp

nic.merit.edu documents/fyi/fyi 18.txt

Internet Wiretap ftp

wiretap.spies.com About/FEATURES

Internet Wiretap Gopher gopher

wiretap.Spies.COM

Internet Town Hall ftp

town.hall.org

IRC Community ftp ftp.eff.org pub/Publications/CuD/Papers/ele ctropolis

ISEA http www.uiah.fi

isea/index.html

ITK http

itkwww.kub.nl itk/itkhome.html 2080

Latin ftp netcom.com

pub/amcgee/latin/

LETRS gopher

gopher.indiana.edu 11/letrs/gopher 1067

MLA93 forum ftp

epas.utoronto.ca pub/cch/mla/

MTV http mtv.com

MUSE http

muse.mse.jhu.edu

Net Rights ftp

ftp.american.edu au/brrec.text

Net Ethics ftp

nic.merit.edu documents/rfc/rfc1087.txt

Net Etiquette Guide ftp ftp.sura.net pub/nic/internet.literature/net

iquette.txt

Net Orgs ftp rtfm.mit.edu pub/usenet/news.answers/net-com

munity/orgs-list

Netizen Anthology ftp

wuarchive.wustl.edu doc/misc/acn/netbook

Netizen Paper ftp

wuarchive.wustl.edu doc/misc/acn/papers/netizens.Z

Networking mail

rre-request@weber.ucsd.edu Subject: archive send network

Online Book Initiative ftp

world.std.com obi/README

Oxford Archive ftp black.ox.ac.uk

ota/

Post-Gutenberg Galaxy ftp

infolib.murdoch.edu.au pub/jnl/harnad.jnl

Privacy ftp

rtfm.mit.edu pub/usenet/alt.privacy/

Privacy forum gopher vortex.com

11/privacy

Reasons for NII ftp ftp.cni.org

CNI/documents/farnet/README

Smiley Dictionary ftp

ftp.gsfc.nasa.gov pub/smiley-dictionary

Smileys ftp ftp.uu.net usenet/comp.sources.misc/volume 23/smiley/part01.Z

Togethernet gopher

gopher.together.uvm.edu

Tribe mail listserv@lists.colorado.edu

“info tribe”

Usenet ftp ftp.eff.org pub/Publications/CuD/Papers/lev iathan

Virtual Community ftp ra.msstate.edu pub/docs/words-l/Net-Stuff/slic

e.of.life

Taking/Web http

minnie.cs.su.oz.au writ/start.html

The WELL ftp ftp.eff.org pub/Net info/Cyber/town-on-inte rnet-highway

UWI Cultural Play http zapruder.pds.med.umich.edu

uwi.html

Writer’s Resources ftp rtfm.mit.edu pub/usenet/news.answers/writing

/resources

EDUCATION/ACADEMIA

AskERIC ftp ericir.syr.edu

pub/

AskERIC Gopher gopher

ericir.syr.edu

AskERIC Cows http

eryx.syr.edu COWSHome.html

Bell-TCI ftp ba.com

pub/BASIC ED CONNECTION/

CAF FTP Archive ftp ftp.eff.org

pub/caf

CAF Web Page http www.eff.org

CAF/cafhome.html

CALICO mail

CALICO@Dukemvs.ac.duke.edu

CAUSE gopher gopher

cause-gopher.Colorado.edu

CELIA ftp archive.umich.edu

celia-ftp/

CoVis http

www.covis.nwu.edu

CWIS List ftp

sunsite.unc.edu pub/docs/about-the-net/cwis/cwi s-l

CWIS Paper ftp sunsite.unc.edu pub/docs/about-the-net/cwis/hal

lman.txt

Cyberion City telnet michael.ai.mit.edu

‘guest’

Distance Ed ftp

una.hh.lib.umich.edu inetdirsstacks/disted:ellsworth

Distance Ed DB mail

n.ismail@open.ac.uk

Distance Ed DB telnet acsvax.open.ac.uk

‘icdl’

Diversity U telnet erau.db.erau.edu

8888

Education Gopher gopher

sci-ed.fit.edu

Educator’s Email ftp

nic.umass.edu pub/ednet/educatrs.lst

Educator’s USENET ftp

nic.umass.edu pub/ednet/edusenet.gde

Empire Schoolhouse telnet nysernet.org

‘empire’

Electronic Academical Village http

jefferson.village.virginia.edu iath/iath pamphlet.html

Exploratorium http

www.exploratorium.edu

High school/Internet ftp

sci-ed.fit.edu pub/Internet/study/

IAT mail INFO.IAT@mhs.unc.edu

“send help”

IAT archive ftp

gandalf.iat.unc.edu user/home/anonftp/guides/

IKE gopher

ike.engr.washington.edu

IKE Web http ike.engr.washington.edu

ike.html

Incomplete Guide K12 ftp ftp.ncsa.uiuc.edu Education/Education Resources/I

ncomplete Guide/

Internet and Ed ftp ftp.msu.edu

pub/education/

IRD/Educators ftp tcet.unt.edu pub/user-

supported/horsehorse/K -12NREN1.1 pub/telecomputing-in fo/IRD

JASON telnet

topcat.bsc.mass.edu

KidLink gopher

kids.ccit.duq.edu

K-12 Briarwood gopher

gopher.briarwood.com

K-12 Essays gopher quest.arc.nasa.gov

00/essay

K-12 NASA http

k12mac.larc.nasa.gov hpcck12home.html

K-12 NREN ftp ftp.u.washington.edu pub/user-supported/horsehorse/K

-12NREN1.1

Learning V mail

tmg@nptn.org

Marcopia http

hakatai.mcli.dist.maricopa.edu

MEU BBS telnet bbs.meu.edu

Networking mail

comserve@vm.its.rpi.edu “send Profess Network”

NPTN ftp nptn.org

pub/info.nptn/basic.guide.txt

Novalink’s Education collectio n http

alpha.acast.nova.edu education.html

OISE Gopher gopher porpoise.oise.on.ca

70

Online LC gopher

pringle.mta.ca

Primary/Sec ftp

nic.merit.edu documents/fyi/fyi 22.txt

Scholarly Communication ftp

borg.lib.vt.edu pub/vpiej-l/reports

Scholary Comm Project http

borg.lib.vt.edu z-borg/www/scholar.html 80

Scholarly Comm/Libraries ftp ftp.cni.org

ARL/mellon/

Scholarly Publishing ftp sunsite.unc.edu pub/docs/about-the-net/trln-cop

yright-paper

Scholarly Societies gopher watserv2.uwaterloo.ca 11/servers/campus/scholars SME mail jbharris@tenet.edu “Please send Subject Matter Ex perts info”

Tech/Schools ftp ftp.u.washington.edu pub/user-supported/horsehorse/r

efuse 1.4

US Dept of Ed gopher

gopher.ed.gov

USENET University ftp nic.funet.fi

pub/doc/uu/FAQ

Willow http shebute.com

Projects/Willow/Willow.HTML

CMC/MEDIA/TECHNOLOGY PROGRAMS

Annenberg http www-penninfo.upenn.edu 1962/penninfo-srv.upenn.edu/900

0/11817.html

CCCC http it.njit.edu njIT/Department/CCCC/default.ht ml

Michigan Tech gopher gilligan.hu.mtu.edu

11

Syracuse gopher

cwis.syr.edu 1/othergophers/ist

UC-Irvine http

www.ics.uci.edu ICShome.html

GOVERNMENT/PUBLIC POLICY

ACE gopher cyfer.esusda.gov

11/ace

Clinton Information ftp

ftp.cpsr.org cpsr/clinton/

Clipper ftp ftp.eff.org

pub/EFF/Policy/Clipper/

Copyright FAQ ftp rtfm.mit.edu pub/usenet/news.answers/law/Cop

yright-FAQ/

FCC ftp ftp.fcc.gov

Federal Info Resources ftp

nic.merit.edu omb/INDEX.omb

Federal Information ftp ftp.nwnet.net user-docs/government/keller-gov

-guide.txt

Fedworld telnet

fedworld.doc.gov

Fedworld Feasibility ftp

ftp.nwnet.net user-docs/government/fedline.tx t

GILS ftp ftp.cni.org

pub/docs/gils/

Govt/Citizenship Information gopher

eryx.syr.edu

Govt Agencies ftp is.internic.net infosource/internet-info-for-ev

erybody/government-agencies

Govt Gophers gopher peg.cwis.uci.edu 11/gopher.welcome/peg/GOPHERS/g ov 7000

Govt Information ftp ftp.nwnet.net user-docs/government/gumprecht-

guide.txt

Economic BB telnet ebb.stat-usa.gov

‘trial’

Electronic Records ftp

ftp.cu.nih.gov nara electronic/

Humanities Intitiative mail

chhenry@vassar.edu

Info Policy mail

listserv@uhupvm1.uh.edu “GET GOODYEAR PRV4N6”

Internet Accounting ftp

ftp.sdsc.edu pub/sdsc/anr/papers/accting.sg. ps.Z

Internet Economics ftp gopher.econ.lsa.umich.edu pub/Papers/Economics of Interne t.txt

Internet Economics Collection http

gopher.econ.lsa.umich.edu EconInternet.html

Internet Policy ftp

nic.merit.edu pub/documents/rfc/rfc1527.txt

Internet Pricing ftp

gopher.econ.lsa.umich.edu pub/Papers/Pricing the Internet .txt

Internet Public Subsidy ftp

ssugopher.sonoma.edu pub/schickele.txt

Library of Congress telnet marvel.loc.gov

‘marvel’

NII FTP ftp

ftp.ntia.doc.gov pub/niiagenda.asc

NII BBS telnet iitf.doc.gov

‘gopher’

NII Gopher gopher

iitf.doc.gov

NII Dialup phone 1-202-501-1920

Open Platform ftp ftp.eff.org

pub/EFF/papers/Open Platform/

Politics gopher fir.cic.net 11/Politics

Tap Info ftp ftp.cpsr.org

%CHECKED

Telecom comments ftp ftp.govt.washington.edu

wutc/

Telecom legislation ftp

ftp.govt.washington.edu legislation.telecom/

US Fed Govt ftp nevada.edu liaison/

US Senate gopher gopher.senate.gov 70

3 INFORMATION SERVICES/ELECTRONIC PUBLICATIONS 19

3 INFORMATION SERVICES/ELECTRONIC PUBLICA-TIONS

This section lists sources of information devoted to the study of CMC, the Internet, or computer network technology.

INFORMATION SERVERS

ALMANAC mail almanac@oes.orst.edu

“send guide”

CCITT Doc Retriev mail teledoc@itu.arcom.ch

“help”

Comserve mail

comserve@vm.its.rpi.edu “send Comserve Helpfile”

Coombs http

coombs.anu.edu.au CoombsHome.html

Concise http info.cern.ch hypertext/DataSources/CONCISE/U

serGuide/Overview.html 80

HCIBIB mail hcibib@bellcore.com

“query:”

RFCs via mail mail

rfc-info@isi.edu “help: ways to get rfcs”

ISDN telnet isdn@bbs.combinet.com

‘isdn’

ISYS mail

comserve@vm.its.rpi.edu “help topics isys”

Virtual Work BBS telnet nis.calstate.edu ‘lewisnts’

DISCUSSION LISTS

AUDIOGRAPHICS mail LISTSERV@cln.etc.bc.ca “Sub Audiographics-L YOUR NAME

”

CBNVEE mail listserv@mcmuse.mc.maricopa.edu

“sub CBNVEE YOUR NAME”

CMC mail

comserve@vm.its.rpi.edu “join CMC YOUR NAME”

Cypherpunks ftp

soda.berkeley.edu pub/cypherpunks/

Cypherwonks mail

Majordomo@Lists.EUnet.fi “sub cypherwonks YOUR NAME”

EduTel mail

comserve@vm.its.rpi.edu “join edutel YOUR NAME”

EDTECH mail

listserv@msu.bitnet “subscribe EDTECH YOUR NAME”

Interpedia mail interpediarequest@telerama.lm. com

“subscribe  YOUR  EMAIL

ADDRESS”

IAMCRNet mail

comserve@vm.its.rpi.edu “join IAMCRNet YOUR NAME”

CNI-DIRECTORIES mail listproc@cni.org “sub CNI-DIRECTORIES YOUR

NAME ”

FUTUREC mail

futurec@uafsysb.uark.edu “sub FUTUREC YOUR NAME”

GOPHERJEWELS mail

Listproc@EINet.net  “SUBSCRIBE  GOPHERJEWELS

YOUR N AME”

HELP-NET mail

listserv@templevm.bitnet “sub HELP-NET YOUR NAME”

HYPEREDU mail

LISTSERV@ITOCSIVM.CSI.IT “SUB HYPEREDU YOUR NAME”

HYTEL-L mail

listserv@kentvm.kent.edu “sub HYTEL-L YOUR NAME”

IECC http

www.stolaf.edu network/iecc.html

IECC Archive ftp

ftp.stolaf.edu pub/iecc/archive.discussion/

InterNIC Announce mail

listserv@is.internic.net “sub announce YOUR NAME”

InterNIC Net Resources mail

listserv@is.internic.net “sub net-resources YOUR NAME”

InterNIC NICS mail

listserv@is.internic.net “sub YOUR EMAIL ADR nics”

InterNIC Net Happenings mail

listserv@is.internic.net “sub net-happenings YOUR NAME”

Interpedia mail interpediarequest@telerama.lm. com

“sub YOUR EMAIL ADDRESS”

IPCT-L mail

listserv@guvm.georgetown.edu “sub IPCT-L YOUR NAME”

IPCT Gopher gopher

guvm.ccf.georgetown.edu

IRVC-L mail

listserv@byrd.mu.wvnet.edu “sub IRVC-L YOUR NAME”

LM NET mail

listserv@suvm.syr.edu “sub LM NET YOUR NAME”

MEDIA-L mail

listserv@bingvmb.bitnet “subscribe MEDIA-L YOUR NAME”

MediaMOO telnet purple-crayon.media.mit.edu

8888

MBU mail

listserv@ttuvm1.ttu.edu “sub MBU-L YOUR NAME”

NETTRAIN mail

listserv@ubvm.cc.buffalo.edu “sub NETTRAIN YOUR NAME”

Net Happenings Archive telnet gopher.cni.org

‘brsuser’

NEW-LIST mail

listserv@ndsuvm1.bitnet “sub NEW-LIST YOUR NAME”

NewJour-L mail

listserv@e-math.ams.org “sub NewJour-L YOUR NAME”

NewNIR-L mail

listserv@itocsivm.csi.it “sub NEWNIR-L YOUR NAME”

NII Regional mail

majordomo@rain.org “subscribe niiregional-l”

NIR mail

mailbase@mailbase.ac.uk “sub NIR YOUR NAME”

OCLC-JOURNALS mail listserv@oclc.org “sub OCLC-JOURNALS YOUR

NAME”

Online-news mail Majordomo@marketplace.com “subscribe online-news YOUR BI O

STATEMENT”

UNITE mail

mailbase@mailbase.ac.uk “join UNITE YOUR NAME”

RESODLAA mail

Listserver@usq.edu.au “sub RESODLAA YOUR NAME”

RITIM-L mail

listserv@uriacc.uri.edu “sub RITIM-L YOUR NAME”

Security mail

problems-request@tdr.com

SCHOLAR mail

listserv@cunyvm.cuny.edu “sub SCHOLAR YOUR NAME”

TECHNET mail

listserv@acadvm1.uottawa.ca  “subscribe  TECHNET  YOUR

NAME”

Technomads mail

technomads-request@ucsd.edu “sub technomads YOUR NAME”

TEI ftp sgml1.ex.ac.uk

tei/

ThesisNet mail thesisnetrequest@seas.upenn.ed u

“sub THESISNET YOUR NAME”

TIPSHEET mail

listserv@wsuvm1.csc.wsu.edu “subscribe TIPSHEET”

VPIEJ-L mail

listserv@vtvm1.cc.vt.edu “sub VPIEJ-L YOUR NAME”

WWW courseware mail www-courseware-request@eit.com

“sub YOUR EMAIL ADDRESS”

WWW literature mail www-literature-request@eit.com “sub YOUR EMAIL ADDRESS”

JOURNALS/NEWSLETTERS

ACN http www.cc.columbia.edu

hauben/home.html

ALAWON mail

listserv@uicvm.uic.edu “sub ALA-WO YOUR NAME”

Bits and Bytes mail

listserv@acad1.dana.edu “SUBSCRIBE bits-n-bytes”

Bits and Bytes archive ftp ftp.dana.edu

periodic/

Boardwatch phone 1-303-973-6038

Boardwatch email mail

jack.rickard@boardwatch.com

CETH Newsletter mail

ceth@zodiac.rutgers.edu

Chaos Corner mail chaosrequest@pelican.cit.corne ll.edu

“sub YOUR NAME”

Chaos Corner Archive ftp pelican.cit.cornell.edu

pub/

Chaos Digest mail

tk0jut2@mvs.cso.niu.edu “sub YOUR NAME”

Cook Report($) mail

cook@path.net “Please send Cook Report info. ”

Convergence mail

AEnigma@Netcom.com “Please send Convergence FAQ”

CPU mail

listserv@cpsr.org “sub CPSR-CPU YOUR NAME”

CPU Archives ftp

etext.archive.umich.edu pub/CPSR/work/

Cu-Digest mail

tk0jut2@mvs.cso.niu.edu “sub YOUR NAME”

Cu-Digest Archive ftp ftp.eff.org

pub/Publications/CuD/

Current Cites ftp

ftp.lib.berkeley.edu pub/Current.Cites/

DECNEWS mail

listserv@ubvm.cc.buffalo.edu “sub DECNEWS YOUR NAME”

EDUPAGE mail

listserv@bitnic.educom.edu “sub EDUPAGE YOUR NAME”

Effector Online ftp ftp.eff.org

pub/EFF/newsletters/Index

EJC/REC mail

comserve@vm.its.rpi.edu “join EJCREC YOUR NAME”

EPIN mail

epin@access.digex.com “Please send EPIN info.”

EJournal mail

listserv@albany.bitnet “sub EJRNL YOUR NAME”

EJournal Archive ftp

ftp.hanover.edu pub/ejournal

EJVC mail

listserv@kentvm.kent.edu “sub EJVC-L YOUR NAME”

EJVC Archive ftp byrd.mu.wvnet.edu

pub/ejvc/

Fed Comm Law J http

www.law.indiana.edu fclj/fclj.html 80

GNN mail

info@gnn.com “Please send GNN info.”

GNN News http

nearnet.gnn.com gnn/news/home/news.html

HOTT-LIST mail

listserv@ucsd.edu “SUBSCRIBE HOTT-LIST”

HPC SELECT NEWS mail sub@hpcwire.ans.net

“Select 601”

INFOBITS mail

listserv@gibbs.oit.unc.edu “sub INFOBITS YOUR NAME”

InfoCycle mail

jwsmith@io.org

Info Policy Online mail

iiaipo-request@his.com “subscribe YOUR NAME”

INFOSYS mail

listserv@american.edu “sub INFOSYS YOUR NAME”

InterText ftp network.ucsd.edu

intertext/

InterText Web Page http network.ucsd.edu intertext/other formats/HTML/IT

toc.html

JASBITS mail jsquires@cerf.net Subject: JASBITS SUBSCRIBE “YO

UR NAME how you knew”

Netweaver mail

comserve@vm.its.rpi.edu “send Netweave Winter91”

Nets and Comm mail

samsam@vm1.yorku.ca

DISTED mail

listserv@uwavm.bitnet “sub DISTED YOUR NAME”

ConneXions mail

ole@csli.stanford.edu” “Please send ConneXions info.”

IBJ ftp nstn.ns.ca

pub/internet-business-journal/

IBJ Gopher gopher

gopher.fonorola.net 11/Internet%20Business%20Journa l

Internet Business Report http

www.doc.ic.ac.uk gnn/mkt/cmp/ibr.html

Internet Letter http mtv.com commercial resources/internet l etter/index.html

Internet Review gopher

gopher.msen.com

Internet World mail

meckler@jvnc.net “Please send Internet World in fo.”

Legion of Doom Tech J ftp ftp.eff.org

pub/Publications/CuD/LOD/

Link Letter ftp

nic.merit.edu newsletters/linkletter

LITA Newsletter ftp dartcms1.dartmouth.edu

litanews/

Matrix News mail

mids@tic.com “Please send Matrix News info. ”

Matrix News some issues ftp

quake.think.com pub/mids/matrix news

MC Journal mail

listserv@ubvm.cc.buffalo.edu “sub MCJRNL YOUR NAME”

Meta ftp

ftp.netcom.com pub/mlinksva

NETNEWS mail

listserv@ndsuvm1.bitnet “sub NNEWS YOUR NAME”

NETNEWS Archive ftp vm1.nodak.edu nnews Online Access mail

70324.343@compuserve.com

PACSR mail

listserv@uhupvm1.uh.edu “sub PACS-P YOUR NAME”

PACSR Index mail

listserv@uhupvm1.uh.edu “get index pr f=mail”

REACH ftp ucsbuxa.ucsb.edu

hcf/

Red Rock Eater mail

rre-request@weber.ucsd.edu Subject: subscribe YOUR NAME

Surfpunk Technical Journal http www.acns.nwu.edu

surfpunk/

TNS ftp ftp.netcom.com

pub/ror/

TidBITS mail

listserv@ricevm1.rice.edu “sub TIDBITS YOUR NAME”

Voices ftp

etext.archive.umich.edu pub/Zines/Voices/

Voorhees Reports mail

markvoor@phantom.com “Please send Voorhees Reports info.”

The WELL gopher

gopher.well.sf.ca.us

Wired mail infobot@wired.com

“get index”

Wired Gopher gopher

wired.com

Wired Web http wired.com

Wired on Web http

www.ncb.gov.sg wired/WoWWW.html

3W mail

3W@ukartnet.demon.co.uk “Please send 3W info.”

3W Web http marketplace.com 0/3W/3W3Info.html 70

4 SOCIETIES AND ORGANIZATIONS 23

4 SOCIETIES AND ORGANIZATIONS

NON-PROFIT/ACADEMIC

ACM SIGCOMM http

gatekeeper.dec.com pub/doc/sigcomm/ccr/overview.ht ml

ASIS mail

asis@cni.org

CCN mail

ccn@civicnet.org

CCN Archive ftp

world.std.com ftp/amo/civicnet/

CIOS mail

comserve@vm.its.rpi.edu “help topics associates”

CIX ftp cix.org

cix/README

CME mail

cme@access.digex.net Subject: ”CME Sub”

CNI ftp ftp.cni.org CNI/

CNRI gopher

ietf.cnri.Reston.va.us

CoSN mail

cosn@bitnic.bitnet “Please send CoSN info.”

CoSN Gopher gopher

cosn.org

CPSR ftp ftp.cpsr.org

cpsr/

CPSR Gopher gopher

cpsr.org

CREN gopher

info.educom.edu

EARN ftp

ftp.earn.net

ECHO telnet echo.lu ‘echo’

EDUCOM mail

inquiry@educom.edu “Please send EDUCOM info.”

EDUCOM Gopher gopher

educom.edu

EFF ftp ftp.eff.org

pub/EFF/about-eff

EFF Web Page http

www.eff.org

EFF Gopher gopher

gopher.eff.org

ENA mail

comserve@vm.its.rpi.edu “send Netweave Winter91”

FARNET ftp farnet.org

farnet/farnet info/

GNA http

uu-gna.mit.edu uu-gna/index.html 8001

HPCC Office gopher

www.hpcc.gov

IATH http

jefferson.village.virginia.edu

IETF ftp cnri.reston.va.us

ietf/

IGC ftp igc.org README

IRVC ftp

byrd.mu.wvnet.edu pub/estepp/IRVC

ISOC ftp nri.reston.va.us

isoc/

ISOC Gopher gopher

gopher.isoc.org

ISTE mail

ISTE@Oregon.uoregon.edu

ITU gopher info.itu.ch

Kestrel Inst http kestrel.edu

pub/mosaic/kestrel.html

MCNC http www.interpath.net

MCNC.html

NIIT http www.esi.com

niit/niit top.html

NPTN ftp nptn.org

pub/info.nptn/

OSF http www.osf.org

index.html 8001

RARE ftp ftp.rare.nl rare/

RIPE ftp ftp.ripe.net

ripe/

RIPE Web http

www.ripe.net

WAENA mail

jwmilton@waena.portal.com

SEA gopher gopher.panix.com

4 SOCIETIES AND ORGANIZATIONS 24

COMMERCIAL TELECOMMUNICATIONS/CONSULTING/PUBLISHERS

Commercial List http

tns-www.lcs.mit.edu commerce.html

AT+T Web Page http

www.research.att.com

Bell Atlantic ftp ba.com

pub/

Bell Atlantic ftp

ba.com

Book Stacks Unlimited telnet

books.com

Branch http branch.com 1080

Career Connections telnet

career.com

CityScape http

www.cityscape.co.uk

Clarknet mail

all-info@clark.net

Consensus ftp netcom10.netcom.com pub/consensus/www/ConsensusFron

tDoor.html

Cyberspace Development mail

development@marketplace.com

DEC ftp gatekeeper.dec.com

info.html

DCI mail

75300.2002@compuserve.com

D3 Electric Catalog http

www.service.com D3/home.html

EBT mail

info@ebt.com

EIT http www.eit.com

EINet http galaxy.einet.net

galaxy.html

Electric Press http

www.elpress.com homepage.html

EUnet http www.eu.net

GEMS http

www.gems.com

Global Village Comm http

info.globalvillag.com welcome.html

GNN Marketplace http

nearnet.gnn.com gnn/mkt/mkt.intro.html

IIA mail

iia.ipo@his.com

InterMart http

kudzu.cnidr.org InterMart/front door.html 80

Internet Dist Svs http www.service.com

home.html

Internet Mall ftp netcom.com

pub/Guides/

Internet Mall gopher peg.cwis.uci.edu 0/gopher.welcome/peg/INTERNET 7 000

InterNex http

www.internex.com

InterSPACE http

cybercafe.demon.co.uk interspace.html

MailCom ftp

netcom1.netcom.com pub/mailcom/README

MarketBase Gopher gopher

mb.com

Maxwell http inept.scubed.com

8001

MCC http

www.einet.net MCC/MCC.html

MIDS ftp ftp.tic.com tic/TIC

MIDS Gopher gopher

gopher.tic.com

MFJ Task force gopher

bell.com

MFJ ftp site ftp bell.com

NOMA mail

natbbs-request@echonyc.com

Novell http

www.novell.com

O’Reilly and Assoc. telnet ora.com

‘gopher’

O’Reilly Resource Center http

nearnet.gnn.com mkt/ora/center.html

Oslonett http

www.oslonett.no

Oslonett Marketplace http

www.oslonett.no html/ON-market.html

Pacific Bell gopher gopher

gw.pacbell.com

Internet.COM mail

info@internet.com “Please send Internet.com info ”

Internet.COM Gopher gopher

gopher.internet.com

MecklerMedia phone

800-MECKLER

MecklerMedia email mail

meckler@jvnc.net

MecklerMedia gopher gopher r2d2.jvnc.net Publishers%20Online/Meckler%20P

ublishing%20Company

NTT Web Page http www.ntt.jp

index.html

Online Career Center gopher garnet.msen.com

9062

PIPEX http

www.pipex.net

Quadralay http

www.quadralay.com

QMS http

www.qms.com www/QMS1.html

RNS http

mosaic.cmc.com rockwell/about.html 8080

SCO Web Page http www.sco.com

index.html

SLED mail sled@drebes.com “info”

5 NEWSGROUPS 25

Sun Microsystems http

sunsite.unc.edu sun/inform/sun-info.html

UTRC http

utrcwww.utc.com UTRC/General/UTRCGen.html

VTLS Web Page http

www.vtls.com

Wavefront http

wavefront.wti.com

XOR  http  xor.com

5 NEWSGROUPS

Newsgroups are sometimes a rich source of information about the Internet, networks, and CMC issues. This section lists newsgroups in which topics related to networks, the Internet, or CMC are discussed. Watch for (FAQ) = periodic posting of Frequently Asked Questions.

COMMERCIAL TELECOMMUNICATIONS/CONSULTING/PUBLISHERS

Beginners news

news.announce.newusers

Computer Archives news

comp.archives

Computer BBS news

comp.bbs.misc

Cu-Digest news

comp.society.cu-digest

Cyberpunk news

alt.cyberpunk

Email news

comp.mail.misc

FAQs news

news.answers

Gopher news

comp.infosystems.gopher

Groupware (CSCW) news

comp.groupware

Hackers news alt.hackers

ISDN news

comp.dcom.isdn

Information Systems news

comp.infosystems

Info Systems – new news

comp.infosystems.announce

Info sys announce mail

infosys@msu.edu

Info sys admin mail

infosys-request@msu.edu

Internet Access news

alt.internet.access.wanted

Internet BBS news

alt.bbs.internet

Internet Best news

alt.best.of.internet

Internet Culture news

alt.culture.internet

Internet Library news

comp.internet.library

Internet Protocols news

comp.protocols.tcp-ip

Internet Relay Chat news

alt.irc

Internet Services news

alt.internet.services

Internet Talk Radio news

alt.internet.talk-radio

Lists of Information news

news.lists

MUDS news

rec.games.mud.announce

New Newsgroups news

news.announce.newgroups

Newsgroups news news.groups

PostScript Net Maps news

news.lists.ps-maps

Technical Reports news

comp.doc.techreports

Telecomm news

clari.nb.telecom

Telecomm news

comp.dcom.telecom

Undernet news

alt.irc.undernet

USENET Culture news

alt.culture.usenet

USENET Humor news

alt.humor.best-of-usenet

USENET Recovery news

alt.usenet.recovery

USENET University news

alt.uu.future

WAIS news

comp.infosystems.wais

World Wide Web news comp.infosystems.www

6 SELECTED BIBLIOGRAPHY 26

6 SELECTED BIBLIOGRAPHY

This section lists useful information sources. For the sake of compactness, the book list here is short. For longer lists, see the online or paper bibliographies available. I’ve prepared a list of many books and article readings in CMC and computer networking. That file is the CMC December bibliography, listed below.

ONLINE BIBLIOGRAPHIES/SYLLABI

Bib List gopher ukoln.bath.ac.uk 11/Link/Tree/Computing/Literatu re 7070

CMC December ftp ftp.rpi.edu pub/communications/internet-cmc .bib

CMC Hacker mail

comserve@vm.its.rpi.edu “get compumed biblio”

CMC Parker mail

listserv@cmuvm.csv.cmich.edu “send netstart info”

CMC Shedletsky ftp tcet.unt.edu pub/telecomputing-info/cmc-bibl

iography.txt

CMC mail

comserve@vm.its.rpi.edu “get compunet biblio”

CMC course/Harrison http www.rpi.edu Internet/Guides/decemj/course/c mc.html

CMC course/Stephen mail

comserve@vm.its.rpi.edu “get compumed syllabus”

CMC ED Higgins mail

comserve@vm.its.rpi.edu “get rescmc syllabus”

CMC ED Paulsen mail

comserve@vm.its.rpi.edu “get disted2 biblio”

CMC ED Rice mail

comserve@vm.its.rpi.edu “get cmced biblio”

CMC Study Review ftp tcet.unt.edu pub/telecomputing-info/cmc-stud

y-review.txt

CSCW Bibliography ftp

ftp.cpsc.ucalgary.ca pub/CSCWbibliography/README

Ed Tech/England mail

comserve@vm.its.rpi.edu “get EdTel biblio”

Electronic Serials ftp

infolib.murdoch.edu.au pub/bib/bailey.bib

Email 1/Thelen mail

comserve@vm.its.rpi.edu “get email 1 biblio”

Email 2/Thelen mail

comserve@vm.its.rpi.edu “get email 2 biblio”

FYI Bibliography ftp

nic.merit.edu documents/fyi/fyi 03.txt

IAT Internet Bib ftp

gandalf.iat.unc.edu user/home/anonftp/guides/irg-14 .txt

Info Nets/Newby mail

comserve@vm.its.rpi.edu “get InfoNet Syllabus”

Information Soc ftp ftp.eff.org

pub/Publications/CuD/CuD/cud5.6 8

Internet Bks Review mail listserv@ubvm.cc.buffalo.edu “GET NETTRAIN REVIEWS NET-TRAIN ”

Internet Bks Quarterman ftp

nic.merit.edu documents/rfc/rfc1432.txt

Internet/NREN Pugh mail

comserve@vm.its.rpi.edu “get internet biblio”

Intro Internetworking ftp

nic.merit.edu documents/fyi/fyi 19.txt

Network Reading List ftp

ftp.utexas.edu pub/netinfo/reading-list/net-re ad.txt

Networked Info Bib ftp

infolib.murdoch.edu.au pub/bib/stanton.bib

New User’s Bib mail

listserv@ubvm.cc.buffalo.edu  “GET  NEWUSERS  FAQ

NETTRAIN”

New User’s Bib (ftp) ftp

sluaxa.slu.edu pub/millesjg/newusers.faq

Telecom/Newby mail

comserve@vm.its.rpi.edu “get Telecomm Syllabus”

WAIS Bibliography ftp

infolib.murdoch.edu.au pub/bib/lincoln.bib

WWW Bib http info.cern.ch hypertext/WWW/Bibliography.html

SPECIAL ISSUES/JOURNALS ON NETWORKS/CMC

Communications of the ACM 1991

Collaborative computing issue Volume 34, number 12, December.

Communication Research 1987 Special issue on Theories of Organizations and Information Technology

Volume 14, number 5, October.

Communication Yearbook 1989 Issues and Concepts in Research on Computer-Mediated Communication Systems

Volume 12, chapter 8.

EJC/REC V3N2 mail

comserve@vm.its.rpi.edu “send EJCTOC V3N293”

Internet Business Journal 1993 Special Issue on Internet Advertising Volume 1, numbers 3/4, September/October 1993.

Journal of Communication 1989 Issue on computer communication affecting social power distribution

Volume 39, number 3.

Online Access 1993 Issue on the Internet

October.

The Serials Librarian 1991 Issue on technology and information for serials

Volume 21, numbers 2/3.

Serials Review 1992 Issue on Economic Models for Networked Information

Volume 18, numbers 1-2.

Scientific American 1991 Issue on computer networks Volume 265, number 3, September

1991.

The South Atlantic Quarterly 1993 Flame Wars: The Discourse of Cyberculture

Volume 92, number 4, Fall 1993.

Wall Street Journal 1993 A Report on Technology Volume CCXXII, number 96, Novem-

ber 15, 1993. Section R.

Written Communication 1991 Issue on computers, language, and writing Volume 8, number 1, January.

BOOKS: a selected listing of particularly useful books.

BIBLIOGRAPHIES

The American Center for the Study of Distance Education 1992 Computer-mediated communication for distance education University Park, PA: The Pennsylvania State University.

Burge, Elizabeth 1992 Computer Mediated Communication and Education: A Selected Bibliography Toronto, Ontario: Distance Learning Office, Ontario Institute for Studies in Education, 252 Bloor St W, Toronto, Ontario M5S 1V6.

Romiszowski, A. J.

1992 Computer-mediated communication: a selected bibliography Englewood Cliffs, NJ: Educational Technology Publications.

THE INTERNET

Carroll, Jim and Rick Broadhead.

1994

The  Canadian  Internet

Handbook

Canada: Prentice-Hall.

Cronin, Mary J.

1993 Doing business on the Internet: how the electronic highway is transforming American companies

New York: Van Nostrand Reinhold.

Braun, Eric.

1994

The Internet Directory New York: Fawcett Columbine.

Dern, Daniel P.

1994 The Internet Guide for New Users

New York: McGraw-Hill.

Estrada, Susan 1993 Connecting to the Internet: An O’Reilly Buyer’s Guide Sebastopol, CA: O’Reilly and Associates.

Falk, Bennett.

1994

The Internet Roadmap San Francisco: Sybex.

Fisher, Sharon 1993

Riding the Internet Highway Carmel, IN: New Riders.

Glister, Paul A.

1993 The Internet navigator: a new user’s guide to network exploration

New York: Wiley.

Hahn, Harley and Rick Stout 1994

The  Internet  Complete

Reference

New York: Osborne McGraw-Hill.

Hahn, Harley and Rick Stout 1994

The Internet Yellow Pages New York: Osborne McGraw-Hill.

Hahn, Harley 1993

A Student’s Guide to Unix New York: McGraw-Hill.

Hardie, Edward T. L. and Vivian Neou 1993

Internet: mailing lists Englewood Cliffs, NJ: Prentice Hall.

Kehoe, Brendan 1994 Zen and the Art of the Internet: A Beginner’s Guide 3rd ed. Englewood Cliffs, NJ: Prentice Hall.

Krol, Ed 1992 The Whole Internet User’s Guide and Catalog Sebastopol, CA: O’Reilly and Associates.

Lambert, Steve and Howe, Walt.

1994 Internet Basics: Your Online Access to the Global Electronic Superhighway

New York: Random House.

Lane, Elizabeth S. and Craig Summerhill 1992 An Internet primer for information professionals: a basic guide to Internet networking technology

Westport, CT: Meckler.

LaQuey, Tracy with Jeanne C. Ryer. Foreword by Al Gore.

1992 The Internet companion: a beginner’s guide to global networking

Reading, MA: Addison-Wesley.

Lynch, Daniel C. and Marshall T. Rose., Eds.

1993

Internet System Handbook Reading, MA: Addison-Wesley.

Malamud, Carl.

1993 Exploring the Internet: a Technical Travelogue

Englewood Cliffs, NJ: Prentice Hall.

Rose, Marshall T.

1993 The Internet Message: Closing the Book with Electronic Mail Englewood Cliffs, NJ: PTR Prentice Hall.

Smith, Richard and Mark Gibbs 1993

Navigating the Internet SAMS Publishing.

Tennant, Roy, John Ober, and Anne G. Lipow 1993 Crossing the Internet Threshold: An Instructional Handbook

Berkeley, CA: Library Solutions Press.

Tolhurst, William A., Mary Ann Pike and Keith A. Blanton 1994

Using the Internet Indianapolis, IN: Que.

Quarterman, John 1994

The  Internet  connec-

tion: system connectivity and configuration Reading, MA: Addison-Wesley.

COMPUTER NETWORKS

Aboba, Bernard.

1994

The  Online  User’s

Encyclopedia

Reading, MA: Addison-Wesley.

Adams, Rick and Frey, Donnalyn.

1993 !%@:: A Directory of Mail Addressing and Networks, 3rd Ed Sebastopol, CA: O’Reilly and Associates.

Kessler, Gary C.

1990 ISDN: concepts, facilities, and services

New York: McGraw-Hill.

LaQuey, Tracy L., ed.

1990 The User’s Directory of Computer Networks

Bedford, MA: Digital Press.

Li, Xia and Nancy B. Crane.

1993 Electronic style: a guide to citing electronic information

Westport, CT: Meckler.

McClure, Charles R., Ann P. Bishop, Philip Doty, and Howard Rosenbaum.

1991 The National Research and Education Network (NREN): Research and Policy Perspectives

Norwood, NJ: Ablex.

Motorola Codex.

1992 The Basics Book of Information Networking

Reading, MA: Addison-Wesley.

Quarterman, John S.

1990 The Matrix: Computer Networks and Conferencing Systems Worldwide

Bedford, MA: Digital Press.

COMPUTER-MEDIATED COMMUNICATION

Chesebro, James W. and Donald G. Bonsall.

1989 Computer-mediated communication: human relationships in a computerized world Tuscaloosa: University of Alabama Press.

Dunlop, Charles and Rob Kling, eds.

1991 Computerization and Controversy: Value Conflicts and Social Choices

Academic Press.

Dutton, William H, Jay G. Blumler, Kenneth L. Kraaemer, Eds.

1987 Wired Cities: Shaping the Future of Communications

Boston, MA: G.K. Hall and Company.

Gattiker, U. E., ed.

1992 Technology-mediated communication

Berlin; New York: W. de Gruyter.

Gilder, George.

1994 Telecosm

New York: Simon and Schuster.

Harasim, Linda M.

1993 Global networks: computers and international communication

Cambridge, MA: MIT Press.

Harasim, L. M., Hiltz, S. R., Teles, L., and Turoff. M.

1994 Learning networks: a field guide

Cambridge, MA: MIT Press.

Hiltz, Starr Roxanne and Murray Turoff.

1993 The Networked Nation: Human Communication via Computer

Rev. Ed. Cambridge, MA: MIT Press.

Lea, Martin, ed.

1992 Contexts of computer-mediated communication

New York: Harvester Wheatsheaf.

Mason, Robin and Anthony Kaye., eds.

1989 Mindweave: Communication, Computers, and Distance Education

New York: Pergamon Press.

Rapaport, Matthew.

1991 Computer mediated communications : bulletin boards, computer conferencing, electronic mail, and information retrieval

New York: Wiley.

Rheingold, Howard.

1993

The  virtual  community:

homesteading on the electronic frontier

Reading, MA: Addison-Wesley.

Sproull, Lee and Sara Kiesler.

1991 Connections: New Ways of Working in the Networked Organization MIT Press.

7 DESCRIPTION OF ITEMS 30

7 DESCRIPTION OF ITEMS

Item Descriptions

Internet-cmc information

describes the Internet and Computer-Mediated Communication list

Internet-cmc tips describes tips for using the Internet and Computer-Mediated Communication list Gold in Networks! a description of gold nuggets in the network, by J. Martin Hitchikers Guide describes the Internet (circa September 1989), by Ed Krol Internet Index The Internet Index [Inspired by ‘Harper’s Index’], by Win Treese New User’s Questions answers questions commonly asked by new users, by Malkin and Marine Surfing the Internet a narrative of what the Internet has to offer, by Jean Armour Polly What is the Internet? by Krol and Hoffman Internet Services FAQ The Frequently Asked Questions (FAQ) list for alt.internet.services Usenet newsgroup AARNet Guide Australian Network Sites and Resources Guide AARNet User Guide Australian Network User’s Guide Big Dummy’s Guide written by Adam Gaffin for the Electronic Frontier Foundation Big Dummy Web Big Dummy’s Guide in HTML Big Dummy Search search the Big Dummy Guide CERFNet Guide by the California Education and Research Federation Network Desktop Internet a DOS hypertext reference to the Internet. De Presno Gopher giving information about Odd de Presno’s shareware book, The Online World De Presno Guide a shareware book, ‘The Online World’ (long) DDN New User Guide Defense Data Network guide for new users InfoPop a WINHELP (hypertext) guide to the Internet, Compuserve, BBS systems, and more, by Wally Grotophorst

Internet Companion(parts)

A Beginner’s Guide to Global Networking by Tracey LaQuey

Internet Guide A guide to Internet/Bitnet, by Dana Noonan Meng’s

Holistic Guide to the Internet

Neophyte Network Knowledge for the Neophyte, Stuff You Need to Know in Order to Navigate the Electronic Village, by Martin Raish

NSF Resource Guide National Science Foundation Guide to Internet Resources NWNet Internet Guide

NorthWestNet’s User Service Internet Resource Guide

NYSERNet Internet Guide

New York State Education and Research Network Guide to the Internet

SURANet Internet Guide

updated as infoguide.MM-YR.txt

SURFnet Guide SURFnet is the Dutch academic network Zen/Art of Internet Zen and the Art of the Internet, by Brendan Kehoe Agricultural Guide Not Just Cows, by Wilfred Drew Electric Mystics Guide A Complete Directory of Networked Electronic Documents, Online Conferences, Se-

rials, Software, and Archives Relevant to Religious Studies, by Michael Strangelove

Library Resources Strategies for Selection and use, Ed. Laine Farley BUBL

The Bulletin Board for Libraries

DOS Internet Kit public domain programs that enable Ethernet or serially-connected PCs to access

Internet services, by Dean Pentcheff

Internet Beginner The Beginner’s Guide to the Internet, a DOS-based tutorial, by Pat Suarez Internet Cruise A Cruise of the Internet, computer-based tutorial for Internet navigators, DOS or

Mac formats

Internet Explorer’s Kit

a hypertext system about the Internet for DOS, by Ernest Perez

Internet Hunt a game for learning about Internet resources Internet Hunt Gopher Hunt Questions, Results, and Comments Internet Navigating Navigating the Internet Workshop List, by Richard J. Smith Internet Mining workshops and electronic event Internet Safari A Classroom User’s Guide to the Internet, by Yvonne Marie Andres Competencies Competencies for Electronic Information Services, by John Corbin Discussion help The art of getting help–guidelines for seeking help in discussion groups Gopherin

a course for cyberspace

HTML Guide A Beginner’s Guide to HTML HTML Guide/Flynn How to write HTML files, by Peter Flynn HTML Tutorial HTML Documents–A Mosaic Tutorial by Wm. Dennis Horn HWGUIDE A hypertext version of Noonan’s Guide to Internet/Bitnet LeJeune Course course materials and files, by Urban A. LeJeune Milles

An Introduction to Using the Internet at St. Louis University School of Law

Monsarrat

Internet Course for Basic or Serious Level

Navigating

an interactive workshop, by Richard J. Smith

Navigating Navigating the Internet, workshop for teaching the Internet via email, by Richard J. Smith

Net workbook by Martin Raish Newcastle/Tyne Network Training Materials and Related Documentation Schneider Bib A bibliography of Internet Training Materials, by Karen Schneider Start

Where to Start for New Internet Users, by James Milles

Trainmat

Training Materials Gopher

Trainpack

repository of training and training-related materials

Wheeler

Internet Training handouts written by Bill Wheeler

Unix Help

Novalinks’s Unixhelp for users

URL

Curling Up to Universal Resource Locators, by Eric S. Theise

URL guide

A Beginner’s Guide to URLs

VAX/VMS Networking

Beginner’s Guide to Networking on the VAX/VMS, by Lucia Ruedenberg

Windows-based Internet training applications for MS-Windows Internet Growth charts showing the Internet’s past and projected growth, by Texas Internet Consulting Internet History Master’s Thesis, Henry Edward Hardy Internet History files about history of the Internet Internet Monthly Report

monthly communications to the Internet Research Group

Internet Statistics NSF Statistics about Internet use Merit

network information center

Merit Gopher network information center NIC/DDN

Network Information Center/Defense Data Network

InterNic Telnet Directory and Database Services Telnet Interface InterNIC Ftp ftp archives for InterNIC InternNIC via email

Mail service

InterNIC Gopher information server for Internet Network Information Center FYI

For Your Information-sub-series of RFCs about topics which relate to the Internet

Internet Drafts draft documents to be submitted ultimately to the Internet Activities Board to be as

RFCs

Internet Standards sub-series of notes within the RFC series which document Internet standards RFC Request for Comments-a document series which describes the Internet suite of protocols and related experiments

RFC Web

Internet Request For Comments (RFC) in html

Internet Web Text an index to the Internet’s resources, John December GNN’s Internet Internet information from Global Network Navigator Nova Links

Nova Southeastern University, in Ft. Lauderdale Florida

Stanford

Stanford Medical School Internet Page

NCIH

North Carolina Information Highway

NOSC Page

list of Internet-related sources, by Richard Bocker

Yale

General Overview of Worldwide Internet Resources

ACM Gopher Association of Computing Machinery’s information server ALAWON archive files from American Library Association CMD

Center for Human-Computer studies

CNI Gopher Coalition for Networked Information, promotes creation of and access to information resources in networked environments

CNI search

Coalition for Networked Information search server

CTR/Columbia Columbia University Center for Telecommunications Research (CTR) Communications Res Cen

Communications Research Centre, Ottawa, Canada

CERN European Laboratory for Particle Physics, Geneva, Switzerland, (and birthplace of the World-Wide Web)

Dern’s Info Daniel Dern’s Internet Info, News and Views EFF Gopher

A gopher information server for the Electronic Frontier Foundation

HelpNet Archives reference base for new users to the Internet HCI ftp site

Human-Computer Interaction (HCI) bibliography repository

HCI Launching Pad Human-Computer Interaction HPCC Office Web Page the National Coordination Office (NCO) for High Performance Computing and Communications

HPCC Gopher the National Coordination Office (NCO) for High Performance Computing and

Communications

HPCWire The High-Performance Computing news and information service, covering workstations through supercomputers

IIASA International Institute for Applied Systems Analysis–Global Environmental Change; Global Economic and Technological Transitions; Systems Methods for the Analysis of Global Issues

ICSI

The International Computer Science Institute

IEEE Gopher Institute of Electrical and Electronics Engineer’s information server IEEE Tech

Portfolio of Emerging Technologies

ISOC Gopher the Internet Society’s Information server ISS

Institute of Systems Science, National University of Singapore

ISRI

Information Science Research Institute, University of Nevada, Las Vegas

Introducing the Internet

Merit’s Directory of miscellaneous introductory Internet information

Library of Congress a directory of Internet User Guides LIT

Laboratoire d’Informatique Technique

Mem U of Newfoundland

Venturing into the Internet collection

MetaCenter

National MetaCenter for Computational Science and Engineering

Meta Net Gopher support for groups and individuals who use The Meta Network MSU Net and DB Network and Database Resources (CWIS, gopher, electronic info.) MSEN Gopher MSEN offers Usenet and Internet Connectivity NEARnet

New England Academic and Research Network

NIST Web Page National Institute of Standards and Technology WWW Home Page NIST Gopher National Institute of Standards and Technology (USA) Gopher NSF Gopher

National Science Foundation (USA) Gopher

NTTC

National Technology Transfer Center

NYSERnet

New York State Education and Research Network

OPTOLINK

services from The International Society for Optical Engineering (SPIE)

OSS

Open Source Solutions, open sources, methods, and products

PSGnet/RAINet networking in the developing world, low-cost networking tools, computer networking

in general

RICIS

Research Institute for Computing and Information Systems

Singapore DMC Digital Media Centre (DMC) Mosaic Home Page Software Tool and Die The World, online service for dialup access to the Internet SRI’s Network Info SRI International SURAnet

Southeastern Universities Research Association Network

SWITCHinfo

Swiss Network

Telecomm Archives files about telecommunications, from the Usenet group comp.dcom.telecom Telecomm research Center for Telecommunications Research, Columbia University, NY UNC’s Sunsite Univ of North Carolina’s collection USENET Repository collection of FAQ’s and files from Usenet newsgroups USENET Periodic Postings

list of Usenet periodic postings

UUNET archive UUNET Communications Services collection UWO’s NIC

Univ of Western Ontario’s Network Information Centre

Washington and Lee Explore Internet Resources WATERS

Wide Area Technical Report Service

WELL Matrix information about cyberspace networks Cyberpoet Guide cyberpoet’s Guide to virtual culture—a multipart guide to the electronic frontier, by

John Frost

Fishnet

a weekly list of odds and ends of postings and information from the Internet

Freeside FAQ Meng surfs the Internet Gopher Jewels subject-oriented gopher list Gopher Jewels .links format Gopher Jewels Info lists information available about Gopher Jewels Hot/Cool List What’s Hot and Cool on the Web Internet Handbook a list of RFCs by subject and category Information Sources the Internet and Computer-Mediated Communication Internet Meta-Index Internet Resources Meta-Index, by Marc Andreessen Internet Reviews reviews of Internet resources Joel’s List

Joel’s Hierarchical Subject Index

MaasInfo Indexes The index of indexs of Internet online files, by Robert Elton Maas May List

lists telnet sites/services, by CC May

Must Have List List of five ‘must have’ Internet Resources Networking Overview Overview of information available Planet Earth

planet earth home page, a list of things on the Internet, by Richard Bocker

Smith’s BigFun List telnet, ftp, and other sites, by Jeremy Smith ThesisNet FAQ a summary of academic resources pertaining to cyberspace Yale Overview General Overview of Worldwide Internet Resources Yanoff’s Services List Special Internet Connections (Internet Services List), by Scott Yanoff Yanoff List HTML Hyptertext markup language version of Yanoff’s Internet Services list Best of Web

will contain a gallery of best of Web Award Winnters

Internet Browsers Sources for Internet Browsers and Client Software Internet Tools EARN guide to network resource tools, by European Academic Research Network Association Internet Tools/EARN HTML

a Guide to Network Resources Tools

Internet Tools NIR A status report on networked information retrieval tools and groups, by Joint

IETF/RARE/CNI Networked Information Retrieval Working Group

Internet Tools/December HTML

HTML version of a list summarizing Internet tools and forums, by John December

Internet Tools/December

a list summarizing Internet tools and forums, by John December

Internet Systems UNITE

a list of Internet tools and systems, by User Network Interface To Everything

Gopher Jewels Search Search the Gopher Jewels Hyper-Index

search for URL entries that have relations with a set of keywords

JumpStation

referencing the information available on the World Wide Web

NorthStar

search indexes the headers of World Wide Web documents

SUSI

Simplified Universal Search Interface

WWW Worm

finds Web files and places them in a database for searching

WWW Wanderers Wanderers, Spiders and Robots BITFTP

obtain files via anonymous ftp on Bitnet

CARL Systems databases for article access and delivery Correct Time/NBS tells National Bureau of Standards correct time ENewsstand

information provided by U.S.magazine publishers

ENewsstand gopher information provided by U.S.magazine publishers Email FTP

get files at anonymous ftp sites via email.

Email Gopher use a gopher via email Email Gopher use a gopher via email Email WWW obtain a web file (e.g., http) via email; URL = Uniform resource locator Email Usenet post to newsgroup via email FAXNET

FAX’s via email ($)

Finger database Access to a database facility via finger Hytelnet assists users in reaching Internet libraries, Free-nets, CWISs, BBSs, and other information sites by telnet

Internet Marketing by the Internet Company Knowbot

provides a uniform user interface to heterogeneous remote information services

Mail name server gives email information about many email names (e.g. IP address given name)

Softlock

selling freely-copyable computer programs and documents

WorldWindow a service of the Washington University Libraries CERT

Computer Emergency Response Team at Carnegie Mellon Univ.

DoD Security DoD Goal Security Architecture (DGSA) FidoNet Guide The Big Dummy’s Guide to Fidonet, by Michael Schuyler FidoNet News archives of the newsgroup comp.org.fidonet FidoNet Nodes a list of systems with FidoNet FidoNet Gateway How to use the UUCP/Fido-Net Gateway Hypercard TCP/IP Implementing TCP/IP communications with HyperCard INET 93

Proceedings of the INET ’93 conference

International Connect International Connectivity Table, by Larry Landweber Internet Country Codes

FAQ about country codes

Internet Domain Names

Relationship of Telex Answerback Codes to Internet Domains

Internet Root Domain lists of Internet hosts Internet Security GAO report on Computer Security (June 1989) InterNetwork Mail methods of sending mail from one network to another, by John Chew and Scott Yanoff Intro TCP/IP describes the Internet protocols List of FTP Sites list of Internet sites accepting anonymous ftp (December 1991), by Jon Granrose PSGnet/RAINet info networking in developing world, low-cost tools, networking in general NREN Information Merit’s directory of National Research and Education Network information NREN Recompetition contains information about the recompetition of the NSFNET backbone project OneNet

a global network of Macintosh computers

SDSC Appl Net Res Group

activities of the San Diego Supercomputer Center Applied Network Research group.

Sprintlink

FTP site for Sprint’s networking activities

Sprintlink Gopher Sprint’s internetworking activities and networking information USENET Maps Maps of Usenet news feeds/backbones BBS Internet List Zamfield’s Wonderfully Incomplete, Complete Internet BBS List Community Networks information about community networks, by Arthur R. McGee Connecting to Internet

What Connecting Institutions Should Anticipate

DLIST

a list of dedicated line Internet providers, by Susan Estrada

Freenet papers papers about network public access FSLIST

the Forgotten Sites List of public access Internet sites, by Louis Raphael

Global Networks an archive for documents to bring the network to lesser-developed areas of the world Internet Access Guide Access Guide to introducing.the.internet, by Ellen Hoffman Internet Access Individual access to Internet, by James Milles Internet Modem Internet service providers in the US accessible through dial-up connections from a

personal computer, by Genevieve Engel

Internet Providers All Providers Alphabetically Internet Providers Non-US

Non-US Internet Providers

K12 Access

documents of California Education Network (CENet) Technical Planning Committee

K12 Gopher

CICnet’s K-12 Internet gopher

NII

National Information Infrastructure

Network Startup NSF-sponsored Network Startup Resource Center PDIAL

The Public Dialup Internet Access List, by Peter Kaminski

NIXPUB

Public/Open Access UNIX, by Bux Technical Services

Registering on the Net Transition and Modernization of the Internet Registration Service, by S. Williamson Rural Nets

Rural Datafication Project

Rural Datafication gopher information about the rural datafication project– Extending Information Highways for Education, Research, and Economic Development in the Great Lakes States

Rural Datafication Web

bring the power of the Internet to rural and otherwise underserved communities

SEA Community Network

Seattle’s community networks

EUnet European Internet-related service organization PSI

Performance Systems International, dial-up Internet access

PSI

Performance Systems International

Pipline

NYC-based Internet gateway

TMN

The Meta Network

The World

public access Unix

Blacksburg

electronic village

CIAO!

British Columbia, Canada

Cleveland

the Cleveland Free-net

Digital City

de Digitale Stad, Amsterdam, Netherlands

M-Net

Michigan Freenet

Nyx

The Spirit of the Night

NCF

National Capital (Canada) Freenet

NS*CAN

Nova Scotia, Canada

Triangle NC linking the people in the Research Triangle area, including Raleigh, Durham, and Chapel Hill, NC, USA

Victoria BC

in British Columba, Canada

Cataloging Internet Res

explores how to catalog electronic information, by Priscilla Caplan

CHAT

Conversational Hypertext Access Technology, natural-language interface

CHAT

natural language interface to information

CIIR

Center for Intelligent Information Retrieval, University of Massachusetts, Amherst

CNI

Coalition for Networked Information

Doc Center Document Center, a hard copy document delivery service for government and industry specifications and standards

DIMUND

Document Understanding Information and Resources

DIMUND FTP Document Understanding FTP Archives CNI TopNode Project part of CNI’s Directories and Resource Information Services Working Group CNI Gopher

Coalition for Networked Information Gopher

CNIDR Gopher Coalition for Networked Information Discovery and Retrieval Gopher CNIDR Web Page Coalition for Networked Information Discovery and Retrieval Home Page IAFA

Internet Anonymous FTP Archives Working Group

Integrated Information

A Vision of an Integrated Internet Information Service, by Weider and Deutsch

Library Special Library special collections on the internet NIR Archives Networked Information Retieval NIR Gopher

Networked Information Retrieval Gopher

OCLC Research Toward Providing Library Services for CMC PCP Principia Cybernetica Project (PCP)–the computer-supported collaborative development of an evolutionary-systemic philosophy

Resource Transponders

by Weider and Deutsch

Retrieval success successful stories of using the Internet for reference, by Karen Schneider Riddle

Rapid Information Display and Dissemination in a Library Environment

UNITE Archive User Network Interface To Everything URN

Uniform Resource Names, by Weider and Deutsch

URI Archive

transcript of Uniform Resource discussion list

UWI

UnderWorld Industries, nodes that wish to share information with other UWI nodes

Commercial List List of Commercial Services on the Web CSOIRG Home Page The Clearinghouse for Subject-Oriented Internet Resource Guides Dartmouth Merged SIGL

Special Interest Group Lists

Electronic Conferences

Directory of Scholarly Electronic Conferences, by Diane K. Kovacs

Electronic Zines/Jrl An Index to Electronic Journals and Zines Electronic Journals Directory of Electronic Journals and Newsletters, by Michael Strangelove Electronic Magazines an archive of various electronic magazines Electronic Resources lists resources (CWIS, texts, guides, WWW, WAIS, etc.) by University of Waterloo

Library

EZines List summary of electronically-accessible zines, by John Labovitz EZines Web

summary of electronic zines by John Labovitz

EINet Galaxy a guide to world-wide information and services Finding Lists how to find LISTSERV lists Free Databases a catalog databases that are available without payment InterActive Yellow Pages

by Lighthouse Productsion–telephone style yellow pages international directory

Interest Groups List The Lists of Lists, a listing of special interest group mailing lists available on the

Interne

Library Access Script a script to login to online card catalogs Library Guide Accessing on-line Bibliographic Databases, by Billy Barron and Marie-Christine Mahe Library Catalogs Noonan

a list of academic and public library catalogs online, by Dana Noonan

Library Catalogs St Geo

Internet - Accessible Library Catalogs and Databases, Art St. George and Ron Larsen

Online books

On-line books

Online bibs

On-line Bibliographies and Journal Contents

Online journals On-line Journals PAML

Publicly Available Mailing Lists, by Stephanie da Silva

Subject Guides subject-specific Internet Resource Guides collected by UMich Subject Guides UMich collection via gopher Subject Guides UMich’s collection via WWW Technical Reports technical reports archive Virtual Tourist geographical directory of WWW Information Services WAIS sources a brief description of the content of many WAIS sources on the Internet, grouped into

relevant categories

WWW Catalogues A collection of WWW Home pages in many areas, by Oliver McBryan WWW Sites comprehensive list–a long list of WWW sites gathered through an automated program, Webwanderer, by Matthew Gray

Berklee

College of Music

CERL

The CERL Sound Group (U of IL)

Internet Sound various documents and programs having to do with sound Internet Talk Radio by Carl Malamud Internet Multicasting FAQ

FAQ for the Internet Multicasting Service

Internet Multicasting WWW

Home page for the Internet Multicasting Service

Leeds/Music

Leeds University (UK) Department of Music

MIDI

Musical Instrument Digital Interface

Multicast Backbone live audio and video multicast virtual network on top of Internet Mbone FAQ Web Frequently Asked Questions (FAQ) on the Multicast Backbone (MBONE) Say

Text-¿Audio (translate text to sound)

IUMA

Internet Underground Music Archive

ANIMA

Arts Network for Integrated Media Applications

CGU (Manchester, UK)

The Computer Graphics Unit - Research

File formats

graphics and sound file formats

GA Tech Graphics Georgia Institute of Technology’s Graphics, Visualization, + Usability Center HTML guide

a guide to writing documents in hypertext-markup language

HTML guide/Tilton Writing Good HTML, by James ’Eric’ Tilton HTML+ DTD

draft of extension specifications to HTML, by Dave Raggett

Hypermedia/Internet Hypermedia and the Internet, by David Green Meta VE

Meta Virtual Environments

Multimedia Index Multimedia Information Sources, by Simon Gibbs Multimedia Survey A Survey of Distributed Multimedia Research, Standards and Products, by RARE Multimedia Lab BU Multimedia Laboratory at Boston University MIT Media Lab access to Massachusetts Institute of Technology’s Media Lab MIT Meida Lab–how to

How To Do Research In the MIT AI Lab

MIT Telemedia Telemedia, Networks, and Systems Group Scientific Visualization Annotated Scientific Visualization URL Bibliography Sunsite Multimedia multimedia presentations based on SunSITE TX A+M Media Lab Texas A and M Hypermedia Research Lab Video

demonstration of vsbrowser, a video file browser

VR Page/Chris Hand Research, Papers, Archives, Events, User Groups and so on VR Web

Virtual Reality Web Page

VR Archive

Sunsite Virtual Reality archive–papers, information, maintained by David Barberi

VSR

Virtual Shared Reality Project

CBC Home Page Canadian Broadcasting Corporation CBC Gopher

Canadian Broadcasting Corporation gopher

ClariNet

complete electronic newspaper delivered via Net

Cyberspace Report a public affairs radio show aired on KUCI, 88.9 FM in Irvine, California ITR

Internet Talk Radio

Live TV demonstration consists of live video and audio using the vsdemo VuSystem application that uses a live television source on the VuNet

Media BBS

Media BBS List, broadcasting/journalism BBS list, by Mark S. Leff

Movies and TV Nova Link’s Movies and TV information MTV

Adam Curry’s server of music information and news

NBC News

NBC nightly news comments/questions

NPR Weekend Ed US National Public Radio Weekend Edition – Weekend and Sunday ONLY NPR Talk/Nation US National Public Radio Talk of the Nation NPR Science Friday US National Public Radio Science Friday Program PWN

PeaceNet World News Service

Journalism

Journalism List of Internet Resources, by John S. Makulowich

Satellite

The Satellite TV Page

TV Schedules the schedules of some television channels/shows VOA

Voice of America (USA)

CCCC Computerized Conferencing and Communications Center at New Jersey Institute of Technology (NJIT)

Communication Archive Sunsite Communication archive–papers, logs, information, maintained by David Barberi

Communication Archive Web Sunsite Communication archive–papers, logs, information, maintained by David Barberi

Computer Network Conf

A Discussion on Computer Network Conferencing, by D. Reed

Hypermedia/Internet Hypermedia and the Internet Interactive Geo Interactive Geographical Index, a way for people to mark where they are Internet Relay Chat(IRC)

real-time, many-many text discussion divided into channels

Free for All

an experiment in a networked hypermedia group bulletin board

Multiple User Dialogue (MUD)

real-time interaction usually for social role-playing

MUD Lists

The Totally Unofficial List of Internet Muds

MUD Page

Experimental MUD WWW Page, by Russ Smith

LISTSERV managing How to Start and Manage a BITNET LISTSERV Discussion Group, by Kovacs, Mc-

Carty, and Kovacs

LISTSERV searching Quick Start Guide to Searching the CARR-L Archives LISTSERV tips List Management Tips for LISTSERV Postmasters and List Owners, by Lisa M. Covi Campus Net

Creating a Campus Network Infrastructure, by Steve Griffin

Campus Net Bib Creating a Campus Network Infrastructure–Bibliography Email Services a list of serves available via email, by David DeSimone Email Understanding special issue of Univ of IL newsletter describing electronic mail Email 101

describes how to use email as well as other Internet features, by John Goodwin

College Email addresses

information on email addresses at colleges

Finding Email addresses

how do I find someone’s email address?

Mail2Html Mailbox to HTML Translator–translate a mailbox file (generated by sendmail, for example), into hypertext

MetaMail multimedia

MIME

Multipurpose Internet Mail Extensions

MIME overview Internet Multimedia Mail with MIME, Emerging Standards for Interoperability, by Nathaniel S. Borenstein MIME info

MIME Overview, by Mark Grand

Pine email

a Program for Internet News and Email

PGP Mail

Pretty Good Privacy, a public key encryption package

PGP/PEM allow you to communicate in a way which does not allow third parties to read messages, certify that the person who sent the message is really who they claim they are

PEM

Privacy Enhanced Email

RIPEM

Riordan’s Internet Privacy Enhanced Mail

Activism

a collection of various online resources about activism

Anonymity FAQ Anonymity on the Internet African

a collection of various online resources about African issues

APC

Association for Progressive Communications

Art/images

a collection of art and images in several formats (ASCII, TIFF, GIF, JPEG)

Artwork

OTIS project to collect online art

ASCII Art

A FAQ about ASCII art, by Jorn Barger

BABEL 94

a listing of computer-orientated abbreviations and acronyms, by Irving Kind

Bordeaux and Prague illustrated fictions of both an artistic and a theoretical nature CIS-AH

Center for Integrative Studies - Arts and Humanities

CTI

Centre for Textual Studies, Oxford

Code of the Geeks a symbolic notation to express Geekness Coke Machines bsy’s List of Internet Accessible Coke Machines Community Nets Surveys

a collection of Community Computer Network Surveys

Community Nets collection of papers about Community Computer Networks rights and access Community Nets weekly report attempts to follow the changing nature of community based and oriented

activities on the internet

Computer Jargon search

search the jargon file on WWW

Computer Jargon The Jargon File (the definitive compendium of hacker slang)

Computer Underground The Social Organization of the Computer Underground, an MS Thesis by Gordon R. Meyer

Computing Dictionary The Free On-line Dictionary of Computing Computing Dictionary The Free On-line Dictionary of Computing Coombs Papers science and humanities papers, bibliographies, directories, theses abstracts and other

high-grade research material

CPET

Catalogue of Projects in Electronic Text

Culture/Tech a collection of texts exploring relationship of technology and culture CyberCafe

conference at the Institute of Contemporary Art in London

Cyber Papers EFF’s collection of cyberspace-related papers Cyberpunk FAQ from the Usenet newsgroup alt.cyberpunk Cyberspace/Language Studying the Virtual World–Cyberspace for Language Scholars, by John December EStyle

review of the book *Electronic Style* by Xia Li and Nancy Crane, by Peter Graham

Electronic Text archive

copy of archive services on etext.archive.umich.edu

Electronic Text Catalogue of projects in electronic text, Center for Text and Technology ETC

Electronic Text Center – University of Virginia

Electronic Word Democracy, Technology, and the Arts, by Richard A. Lanham English Server examine the possibilities of collaborative, community-run communications FineArt Forum a newsletter about online art Friends + Partners USA and Russia information and communication FreeNet93

international conference on community networking August 18 and 19, 1993, in Ottawa

Future Culture FutureCulture FAQ, or Cyberography, is maintained by Andy Hawks Gender

a collection of various online resources about gender issues

Gender Issues Gender Issues in Computer Networking, by Leslie Regan Shade Hacker Crackdown text of ‘The Hacker Crackdown-Law and Disorder on the Electronic Frontier’ by Bruce

Sterling

Hacker’s Dictionary a searchable index of Hacker’s Jargon Humanities Ad Hoc Steering Committee of the National Initiative for Humanities and Arts Computing

Humanities/computing Research and Education Applications of Computing in Humanities Newsletter Hypertext/rhetoric Hypertext and the Rhetorical Contract, by Wm. Dennis Horn Indigenous

a collection of various online resources about indigenous peoples issues

Internet demographics a survey by Texas Internet Consulting Internet Glossary search a collection of Internet terms Internet Glossary Internet-specific terms defined Internet Wiretap electronic books and information Internet Wiretap Gopher

electronic books and information

ISEA

International symposium on electronic art

ITK Instituut voor Taal- en Kennistechnologie, Institute for Language Technology and Artificial Intelligence

Latin

a collection of various online resources about Latin issues

LETRS

Library Electronic Text Resource Service

MLA93 forum Reconfiguring the Discipline in the Electronic Age MTV

music-oriented information server run by Adam Curry

MUSE John Hopkins University journals available to students and researchers from their networked desktop computers

Net Rights

Bill of Rights and Responsibilities for the Electronic Community of Learners

Net Ethics

Ethics and the Internet

Net Etiquette Guide by Arlene H. Rinaldi Net Orgs Outposts on the Electronic Frontier, International, Groups Supporting the Online Community

Netizen Anthology The Netizens and the Wonderful World of the Net–an anthology, Ronda Hauben,

Micahel Hauben

Netizen Paper by Michael Hauben, The Net and Netizens Networking

Networking on the Network, by Phil Agre

Oxford Archive Oxford Text Archive Post-Gutenberg Galaxy

The Fourth Revolution in the Means of Production of Knowledge, by Steven Harnad

Privacy

archive about network privacy, from usenet group alt.privacy

Privacy forum materials for the Privacy forum, including all Privacy Forum Digest issues and

materials

Reasons for NII Federation of American Research Networks (FARNet) Network Stories for 51 Reasons

to Build the National Information Infrastructure

Smiley Dictionary unofficial list of smilies Smileys

a smiley server by David W. Sanderson

Togethernet

on-line information and communication network for sustainable planetary future

Tribe

an email list for no purpose but for gathering people together

Usenet

Searching for the Leviathan in Usenet, by Rich MacKinnon

Virtual Community A Slice of Life in My Virtual Community, by Howard Rheingold Taking/Web Taking Over the Web, An agenda for re-colonizing the Internet with the general population

The WELL

Small Town on the Internet Highway

UWI Cultural Play UWI’s Web’s Edge/UnderWorld Industries’ Cultural Playground Writer’s Resources Internet Writer Resources, Compiled/Edited/Maintained by L. Detweiler AskERIC

automated information outlet for ERIC system

AskERIC Cows web server for ERIC educational resources Bell-TCI

Bell Atlantic-TCI project

CAF FTP Archive Computers and Academic Freedom Archive CAF Web Page Computers and Academic Freedom Web Page CALICO

Computer Assisted Language Instruction Consortium

CAUSE gopher association for managing and using information technology in higher education CELIA

Computer enhanced language instruction archive

CoVis

Learning Through Collaborative Visualization, Northwestern University

CWIS List

Campus-Wide Info Systems list

CWIS Paper

Campus-Wide Info Systems by Judy Hallman

Cyberion City virtual environment built by schoolchildren Distance Ed DB A database from the International Centre for Distance Learning at the UK Open

University

Distance Ed DB contains full text of Mindweave , edited by Mason and Kaye Diversity U

Diversity University, an experiment in interactive learning through internet

Education Gopher Florida Tech’s collection of educational resources Educator’s Email An Educator’s Guide to E-Mail Lists Educator’s USENET Ednet Guide to Usenet Newsgroups Empire Schoolhouse Empire Internet Schoolhouse, a selection of K-12 resources, projects and discussion

groups

Exploratorium a collage of 650 interactive exhibits in the areas of science, art, and human perception High school/Internet a case study of an east central Florida high school, by Michael Anthony Gallos to the

internet

IAT

Institute for Academic technology (IBM and UNC)

IAT archive

Institute for Academic technology ftp site

IKE

IBM Kiosk for Education

IKE Web

IBM Kiosk for Education

Incomplete Guide K12 Internet and other telecommunications opportunities for k-12 Internet and Ed by Noel Estabrook IRD/Educators Internet Resource Directory for Educators JASON

introduce teachers to the world of telecommunicating

KidLink

gopher aimed at 10-15 year olds

K-12 Briarwood collection of K-12 resources of the Internet, including a Curriculum Database K-12 Essays

Networks-Where Have you Been All My Life-Essay Contest Winners

K-12 NASA

NASA Langley Research Center’s HPCC K-12 Program

K-12 NREN

Equity, Access, and a Trojan Horse

Learning V Learning Village, Internet access to schools, by National Public Telecomputing Network and Ameritech

Marcopia

Center for Learning and Instruction

MEU BBS

Mind Extension University Bulletin Board System

Networking

Networking on the Network, an essay by Phil Agre

NPTN

National Public Telecomputing Network

OISE Gopher Ontario Institute for Studies in Education Online LC

Mount Allison University’s Online Learning Centre–tele-education

Primary/Sec

Answers to Commonly Asked Primary and Secondary School Internet User Questions

Scholarly Communication

Quarterly Technical Reports

Scholary Comm Project

Scholar – WWW

Scholarly Comm/Libraries Published by The Association of Research Libraries for The Andrew W. Mellon Foundation

Scholarly Publishing Model University Policy Regarding Faculty Publication in Scientific and Technical

Scholarly Journals A Background Paper and Review of the Issues

Scholarly Societies Gophers of Scholarly Societies SME

Subject Matter Experts – Opportunity for Online Teaching and Learning

Tech/Schools Technology Refusal and the Organizational Culture of Schools, by Steven Hodas US Dept of Ed United States Department of Education USENET University an online a society of people interested in learning, teaching or tutoring Willow

The Willow Opportunity Center

Annenberg

Annenberg School for Communication at the University of Pennsylvania

CCCC Computerized Conferencing and Communications Center at New Jersey Institute of Technology (NJIT)

Michigan Tech CCLI (Center for Computer-assisted Language Instruction) Syracuse

School of Information Studies

UC-Irvine

Information and Computer Science

ACE

Americans Communicating Electronically

Clinton Information a collection of presidential and government information Clipper

an encryption chip whose key the government will hold

Copyright FAQ includes issues about Usenet and Internet communication FCC

Federal Communications Commission (USA)

Federal Info Resources documents from the Information Policy Branch, Office of Information and Regulatory

Affairs, Office of Management and Budget of the U.S. government

Federal Information Federal Information and the Internet by Anna Keller Fedworld

a system for government information

Fedworld Feasibility a report on government-wide information locator system by NTIS GILS

Government Information Locator Service

Govt Agencies government agencies on the internet Govt Gophers a collection of government gophers Govt Information Internet Sources of Government Information by Blake Gumprecht Economic BB Commerce Department’s Economic Bulletin Board (EBB) Electronic Records U.S. National Archives and Records Administration Humanities Intitiative National Initiative for Humanities and Arts Computing Info Policy

Information Policy for Electronic Information Resources, by Mary Lou Goodyear

Internet Economics Some Economics of the Internet by Jeffrey K. MacKie-Mason Internet Economics Collection a collection of documents that have to do with the economics of the Internet, information goods, and related issues, by Hal Varian

Internet Policy What Should We Plan Given the Dilemma of the Network?, by G. Cook Internet Pricing Pricing the Internet by MacKie-Mason and Varian Internet Public Subsidy

The Economic Case for Public Subsidy of the Internet, by Sandra Schickele

Library of Congress Marvel (Machine-Assisted Realization of the Virtual Electronic Library) NII FTP

National Information Infrastructure

NII BBS

The White House Information Infrastructure Task Force BBS

NII Gopher

The White House Information Infrastructure Task Force BBS

NII Dialup

The White House Information Infrastructure Task Force BBS

Open Platform Public Policy for the Information Age, EFF Politics

a collection of etexts on miscellaneous political topics

Tap Info

reports on activities relating to federal information policy

US Fed Govt

Internet resources for U.S. Federal Government Information, by Maggie Parhamovich

ALMANAC

A service for multi-media document and information delivery and database functions

Comserve

online service for human communication research

Coombs

information for researchers in the field of social sciences and humanities

Concise

database about networks, networking tools and projects

HCIBIB A mail-based retrieval system interface to a database related to Human-Computer Interaction (HCI).

RFCs via mail Request For Comments–documents about various issues for discussion, covering a

broad range of networking issues.

ISDN

the Combinet, Inc. ISDN deployment database

ISYS

A full text searching/retrieval facility of Comserve’s databases or hotline discussions

Virtual Work BBS Virtual Workplace Bulletin Board AUDIOGRAPHICS audiographics in the context of distance education CBNVEE

educational environments using networked virtual reality software

CMC

Computer-Mediated Communication Comserve Hotline

Cypherpunks

mailing list archive–encryption, remailers, etc.

Cypherwonks general cutting-edge ‘cyberspatial development’ projects such as in cryptography, dig-

ital cash, and ‘electronic democracy’

EduTel

Comserve Hotline use of computers in education

EDTECH

Educational Technology

Interpedia

a public-domain Internet encylopedia

IAMCRNet Comserve Hotline International Association for Mass Communication Research Network

CNI-DIRECTORIES Coalition for Networked Information Directories. FUTUREC

FutureCulture e-list

GOPHERJEWELS sharing interesting gopher finds HELP-NET

For people new to the Internet

HYPEREDU HYPERTEXT in EDUCATION, of hypertexts and hypermedia in education– advantages versus disadvantages

HYTEL-L

announcements of new versions of the popular HYTELNET program

IECC

Intercultural Email Classroom Connections Lists

IECC Archive Intercultural Email Classroom Connections Lists InterNIC Announce announcements list for Internet Network Information Center InterNIC Net Resources

resources only, low volume

InterNIC NICS Internet Network Information Center Network Information Centers list InterNIC Net Happenings

conferences, publications, political news, resources, newsletters, etc.

Interpedia

an online Internet encyclopedia

IPCT-L The Interpersonal Computing and Technology List, issues in education, especially networking and computing.

IRVC-L

foster, ecourage, advance, and communicate research on virtual culture

LM NET

A World-Wide Discussion Group For School Library Media People

MEDIA-L

Media in Education

MediaMOO

Netoric–Tuesday 8PM EST, Electronic Conferences, CMC Workshops

MBU

Megabyte University (Computers and Writing)

NETTRAIN a discussion list about training/support of others in using the resources available on Bitnet and Internet

Net Happenings Archive

search net-happenings mailing list archive

NEW-LIST

discusses new LISTSERV mailing lists.

NewJour-L

new electronic journals/newsletters

NewNIR-L announcements of new Network Information Retrieval and Online Public Access Catalogue Services

NII Regional

National Information Infrastructure (USA) Regional Project Development

NIR Network Information Retrieval. List dedicated to studying networked information retrieval tools and processes.

OCLC-JOURNALS OCLC electronic journal publishing Online-news

online newspapers and magazines

UNITE

interface to all network services.ng. theme

RESODLAA

distance education and open learning research

RITIM-L

Research Institute for Telecommunications and Information Marketing discussion list

Security

the public disclosure of bugs, system problems, viruses

SCHOLAR text analysis and natural language processing, funded by the Andrew W. Mellon Foundation

TECHNET

Technical Support for Education and Research

Technomads

wireless, mobile data comm networks

TEI

Text Encoding Initiative, issues relating to text encoding and SGML

ThesisNet criticism of works-in-progress related to all aspects of human interaction with and within cyberspace

VPIEJ-L

Electronic publishing issues, especially Scholarly Electronic Journals

WWW courseware new or existing educational applications of the World-Wide Web WWW literature new or existing literary resources/works on the World-Wide Web ACN

Amateur Computerist Newsletter

ALAWON

American Library Association (ALA) Online Newsletter

Bits and Bytes a weekly electronic newsletter about miscellaneous high-tech issues Boardwatch

Boardwatch Magaine BBS, voice line for information

Boardwatch email Boardwatch Magaine BBS CETH Newsletter Center for Electronic Texts in the Humanities, Princeton and Rutgers Chaos Corner Dr. Chaos’s random and interesting things about computers, networks, and other

things

Chaos Digest

frauds, hacking, swapping, legislation

Cook Report($) a monthly newsletter on NREN, National Information Infrastructure, and Internet

commercialization

Convergence

entertainment in the age of the information super-highway

CPU

electronic journal for computer industry workers

Cu-Digest technical, social, ethical, legalekly electronic journal/newsletter devoted to many computing issues

Current Cites A journal which provides citations and brief annotations for articles from 30 journals

in networks and information and computer technology

DECNEWS from Digital Equipment Corporation’s Education Business Unit for the education and research communities worldwide

EDUPAGE

summary of information technology news

Effector Online Electronic Frontier Foundation Newsletter EJC/REC Electronic Journal of Communication/La Revue Electronique de Communication Covers communication theory, research, practice, and policy

EPIN Electronic Public Information Newsletter EJournal

Concerned with implications of electronic networks and texts

EJournal Archive back issues of EJournal EJVC Electronic Journal on Virtual Culture, communicate scholarly thought on aspects of virtual [CMC] culture.

Fed Comm Law J Federal Communications Law Journal Web Page GNN

Global Network Navigator

HOTT-LIST

(Hot Off The Tree) advances in computer, communications, and electronics technology

HPC SELECT NEWS The Electronic News Magazine for High-Performance Computing INFOBITS electronic service of the Institute for Academic Technology’s Information Resources Group

InfoCycle

technology management

Info Policy Online newsletter of the Information Industry Association INFOSYS

The Electronic Newsletter for Information Systems

InterText

an electronically distributed magazine of fiction

JASBITS

A newsletter of miscellaneous net information, editor James A. Squires

Netweaver

The Newsletter of the Electronic Networking Association

Nets and Comm Networks and Community, a newsletter of FUTURE DATA DISTED

Online Journal of Distance Education and Communication

ConneXions

Newsletter on information on networking

IBJ

Internet Business Journal

IBJ Gopher

Internet Business Journal Gopher

Internet Letter monthly newsletter about Internet Resources Internet Review An ‘irregular and on-line journal of new and interesting stuff on the net’ Internet World a paper journal from MecklerMedia Link Letter

Merit/NSFNET newsletter.

LITA Newsletter Library and Information Technology Association Matrix News

(monthly online newsletter, Covers crossnetwork issues.

MC Journal

The Journal of Academic Media Librarianship

Meta net community interests–cryptograhy, electronic publishing, free software, internet, privacy

NETNEWS

newsletter for network resources by Metronet (St. Paul, MN)

Online Access oriented to BBS systems PACSR

The Public-Access Computer Systems Review

PACSR Index The Public-Access Computer Systems Review Index REACH

Research and Education Applications of Computers in the Humanities Newsletter

Red Rock Eater Red Rock Eater News Service, Phil Agre’s periodic messages about Internet Surfpunk Technical Journal ‘a dangerous multinational hacker zine originating near BARRNET in the fashionable western arm of the northern California matrix’

TNS

Telelearning Network Synthesizer, Edited by Tom Copley

TidBITS a free weekly electronic publication that reports on interesting products and events in the computer industry

Voices

Voices from the Net Magazine

Voorhees Reports court and regulatory developments involving telecommunications, intellectual prop-

erty, and information

The WELL

The Whole Earth Electronic Magazine

Wired

a magazine of cyberspace and technology

3W

World Wide Web Newsletter

3W Web

World Wide Web Newsletter

ACM SIGCOMM Association for Computing Machinery (ACM) Special Interest Group on

Communications

ASIS

American Society for Information Science

CCN The Center for Civic Networking, applying information infrastructure to the broad public good

CIOS Communication Institute for Online Scholarship, a non-profit organization for online communication scholarship

CIX Commercial Internet eXchange Association, non-profit trade association promoting communications services.

CME

Center for Media Education, promotes democratic potential of electronic media.

CNI Coalition for Networked Information, promotes creation of and access to information resources in networked environments

CNRI Corporation for National Research Initiatives, develop information processing technology CoSN Consortium for School Networking, helps educators and students access information and communications resources

CoSN Gopher Consortium for School Networking CPSR Computer Professionals for Social Responsibility, alliance of computer professionals who discuss the impact of computer technology on society.

CPSR Gopher Computer Professionals for Social Responsibility, alliance of computer professionals

who discuss the impact of computer technology on society.

CREN

supports low-cost access to worldwide electronic networking for education and research

EARN

European Academic Research Network

ECHO

European Commission Host Organisation

EDUCOM dedicated to the study and access of information technology in higher education and integrating technology into learning and research

EFF public interest organization to educate public about computer and communication technologies.

EFF Web Page Electronic Frontier Foundation Web Page ENA

Electronic Networking Association, promotes electronic networking

FARNET non-profit corporation to advance the use of computer networks for research and education.

GNA Globewide Network Academy, a non-profit corporation, affiliated with the Usenet University, goal is create an fully accredited online university

HPCC Office the National Coordination Office (NCO) for High Performance Computing and Communications

IATH

Institute for Advanced Technology in the Humanities

IETF

Internet Engineering Task Force

IGC Institute for Global Communications, improving global network communication and information exchange

IRVC

Institute for Research on Virtual culture

ISOC The Internet Society, supports the development of the Internet and promotes education and applications.

ISTE

International Society for Technology in Education

ITU International Telecommunication Union, a United Nations agency which coordinates telecommunications

Kestrel Inst a non-profit computer science research institute methods for incremental automation of the software

MCNC electronic and information technologies, private/non-profit, formerly the Microelectronics Center of North Carolina

NIIT

National Information Infrastructure Testbed

NPTN

National Public Telecomputing Network

OSF

Open Software Foundation

RARE

Reseaux Associes pour la Recherche Europeenne, European networking organizations

RIPE

Reseaux IP Europeen, ensures coordination to operate a pan-European IP network.

RIPE Web

RIPE Network Coordination Centre

WAENA

WideArea Educational Network

SEA

Society for Electronic Access

Commercial List for a longer list of Commercial Services on the Web AT+T Web Page Web server for AT+T Bell Laboratories Bell Atlantic Bell Atlantic FTP site Bell Atlantic Bell Atlantic gopher Book Stacks Unlimited

on-line Bookstore and Readers’ Conference System

Branch

Branch Information Services, advertising and marketing for the Internet

Career Connections recruiting, advertising, and career searching tool CityScape

CityScape Internet Services, Ltd.

Clarknet

World Wide Web publishing and Internet connections

Consensus

software support for collaboration, including groupware

Cyberspace Development

consulting in building an Internet storefront

DEC

Digital Equipment Corporation, Marketing Home Page

DCI Digital Consulting-computer/telecommunications conferences (e.g. Distributed Computing World, etc)

D3 Electric Catalog

pc products

EBT Electronic Book Technologies, Inc., makers of DynaTextT, transforms ISO standard SGML data and other multi-media forms into dynamic electronic books EIT

Enterprise Integration Technologies

EINet

a commercial provider of network communication and information services

Electric Press give your organization its own business presence on the Internet EUnet

Network Information Services

GEMS

Global Electronic Marketing Service

GNN Marketplace Global Network Navigator (GNN) Marketplace IIA Information Industry Association, 500 member companies in the generation, processing, distribution and use of information

InterMart

Mac/PC/Unix software, books, magazines

Internet Dist Svs electronic publishing and distribution of catalogs and product information Internet Mall A monthly list of commercial services available via Internet, by Dave Taylor Internet Mall A monthly list of commercial services available via Internet, by Dave Taylor InterSPACE

Internet Information Publishers

MailCom

a telecommunications publishing and training firm located in Oakland, California.

MarketBase Gopher on-line catalog of goods and services MCC

Microelectronics and Computer Technology Corporation

MIDS Matrix Information and Directory Services, a private consulting partnership based in Austin, Texas.

MFJ Task force an ad-hoc coalition of the seven regional Bell operating companies MFJ ftp site

an ad-hoc coalition of the seven regional Bell operating companies

NOMA

National Online Media Association

O’Reilly and Assoc. publisher of books for UNIX, open systems, and networking. O’Reilly Resource Center

O’Reilly and Associates, Inc.

Oslonett

a Norwegian company providing Internet and Unix services

Internet.COM The Internet Company, consulting in network access and commercial use of the

Internet

MecklerMedia publishers of Internet books and conferences NTT Web Page Nippon Telegraph and Telephone PIPEX

The Public IP Exchange

Quadralay advanced object-oriented UNIX application tools that simplify and speed the software development process

QMS imaging solutions in the area of electronic publishing, advanced imaging, office automation

RNS

Rockwell Network Systems, high-performance, standards-based networking equipment

SCO Web Page The Santa Cruz Operation SLED

Stable Large Email Database, an email and public key directory

Sun Microsystems Archives, Newsgroups and More! UTRC

United Technologies Research Center

VTLS Web Page automation solutions for libraries and other information centers worldwide Wavefront

graphics for animation

XOR

Network Engineering

Beginners

for people new to USENET

Cu-Digest

Computer Underground Digest

Cyberpunk

cyberpunk/cyberspace culture and literature

FAQs

Frequently Asked Questions (and answers)

Info Systems – new (Moderated) announcement of new information systems or resources Info sys announce send announcements to this address Info sys admin send administrivia to this address MUDS

Multiple User Dialogues

Undernet

an alternative to the current EFNet IRC system

Bib List

list of bibliographies related to computing literature

CMC December bibliography of selected readings in CMC and networking, by John December CMC Hacker

CMC bibliography, by Kenneth L. Hacker

CMC

communication

CMC course/Harrison graduate CMC theory, Rensselaer Polytechnic Institute, Teresa Harrison, John De-

cember, Emilie Goulde, Laura Gurak

CMC course/Stephen description of a CMC course CMC Study Review Ed CMC, a field study of recent research Ed Tech/England for a course, Introduction to Educational Telecommunications Email 1/Thelen email in social contexts (A-H)

Email 2/Thelen email in social contexts (I-Z) Info Nets/Newby course outline and bibliography for a course ‘Information Networking’ Information Soc bibliography by Rob Kling Internet Bks Review reviews of Internet books, from NETTRAIN list Internet/NREN Pugh an annotated bibliography of the Internet and NREN by Pugh New User’s Bib a short bibliography for new users, from NETTRAIN list Telecom/Newby course outline and bibliography for a course entitled ‘Telecommunications’ WWW Bib

Bibliography for the World Wide Web

EJC/REC V3N2 Electronic Journal of Communication special issue on CMC